### v3

# A fast particle-field method for 3D Chemotaxis systemsJingyuan Hu, Zhongjian Wang, Jack Xin and Zhiwen Zhang

Abstract
In this paper, we develop a novel numerical framework, namely the stochastic interacting particle-field method with particle-in-cell acceleration (SIPF-PIC), for the efficient simulation of the three-dimensional (3D) parabolic-parabolic Keller-Segel (KS) systems. The SIPF-PIC method integrates Lagrangian particle dynamics with spectral field solvers by leveraging localized particle-grid interpolations and fast Fourier transform (FFT) techniques. For $P$ particles and $H$ Fourier modes per spatial dimension, the SIPF-PIC method achieves a computational complexity of $O(P+H^{3}\log H)$ per time step, a significant improvement over the original SIPF method (proposed in [SIPF1]), which has a computational complexity of $O(PH^{3})$, while preserving numerical accuracy. Moreover, we carry out a rigorous error analysis for the proposed method and establish rigorous corresponding error estimates. Finally, we present numerical experiments to validate the convergence order and demonstrate the computational efficiency of the SIPF-PIC method. To the best of our knowledge, this work represents the first numerical approach that can resolve complex 3D blowup dynamics beyond single-point collapse, including the formation of ring-shaped singular structures.
keywords:
Parabolic-parabolic Keller-Segel (KS) system, stochastic interacting particle-field (SIPF) algorithm, particle-in-cell (PIC), three-dimensional (3D) simulations, finite-time blowup, convergence analysis.
{MSCcodes}
35K55, 65M12, 65M70, 65M75, 65Y20.
1 Introduction
The Keller-Segel (KS) system [KS] describes chemotaxis-
driven aggregation processes, particularly modeling the dynamics of the slime mold amoeba Dictyostelium discoideum in response to the gradients of the chemoattractant. Here, we study the parabolic-parabolic KS system on a bounded spatial domain $\Omega\subset\mathbb{R}^{d}$:
(1) $\left\{\begin{aligned} \partial_{t}\rho&=\nabla\cdot\left(\mu\nabla\rho-\chi\rho\nabla c\right),\\ \epsilon\partial_{t}c&=\Delta c-k^{2}c+\rho,\quad\mathbf{x}\in\Omega,\ t\in[0,T],\end{aligned}\right$.
where $\rho(\mathbf{x},t)$ is the density of the particles, $c(\mathbf{x},t)$ is the concentration of the chemoattractant, and $\chi,\mu,\epsilon,k\,>0$ are physical constants. Formally taking $\epsilon=0$ yields a parabolic-elliptic KS system that models instantaneous chemical equilibration.
The KS system demonstrates interdisciplinary utility, with applications spanning biological, ecological, and medical domains. In biology, these models describe cell aggregation and migration in bacterial colonies and cancer cells driven by chemotaxis [Transport]. In ecology, they analyze population dynamics shaped by chemotactic strategies, such as algal bloom dynamics and resource-seeking behaviors [crossref2]. In medical research, the KS system offers insights into both physiological mechanisms (e.g., tissue repair and immune response) and pathological processes (e.g., cancer metastasis) [crossref3].
Due to the nonlinear and potentially singular behavior of KS systems, especially in the presence of blow-up phenomena [nagai1995blow, herrero1998self], numerical methods have become essential for analyzing their solutions. Mesh-based numerical methods are among the most widely used approaches for solving the KS system. Chertock et al. [pos_preserving] proposed a high-order hybrid finite volume and finite difference scheme with positivity-preserving properties for two-dimensional (2D) problems. Shen et al. [shen2020unconditionally] proposed an energy dissipation and bound-preserving scheme that is not restricted to specific spatial discretizations. Chen et al. [chen2022error] developed a fully-discrete finite element scheme for the 2D parabolic-elliptic case, and demonstrated that the scheme exhibits finite-time blow-up under assumptions analogous to those of the continuous case. Other works include a positivity-preserving and asymptotic-preserving semi-discrete scheme [LiuJian-Guo2018Paap], and a saturation concentration setting [hillen2001global] which prevents blow-up and yields spiky solutions. Beyond these 2D cases, Gnanasekaran et al. [3Dblowup] simulated a 3D blow-up solution in a cube with a finite difference scheme in a short time.
In addition to the Eulerian discretization methods mentioned above, there have been notable advancements in the Lagrangian framework for the KS system (1) and related equations. The parabolic-elliptic case is comparatively simpler, characterized by a memoryless particle interaction system [velazquez2004point] with singularities arising when a pair of particles approach closely. Several particle methods seek to circumvent the singularity; Havskovéč and Ševčovič [havskovec2009stochastic] employed a regularized interaction potential, and Liu et al. [liu2017random] developed a random particle blob method with a mollified kernel. Convergence studies for these approaches have been conducted in [havskovec2011convergence, mischler2013kac] and [liu2019propagation], with mean-field equations serving as a key tool [mischler2013kac]. Other works include the computation of 2D advective parabolic-elliptic KS systems with passive flow [khan2015global], and the reduction of computational costs via the random batch method for particle interactions [random_batch].
Existing Lagrangian frameworks for parabolic-elliptic cases cannot be directly extended to fully parabolic systems, as the chemo-attractant concentration no longer achieves rapid local equilibration and the drift terms depend on historical particle positions [chen2022mean, fournier2023particle]. Stevens [stevens2000] developed an $N$-particle system for the fully parabolic case, representing the chemo-attractant concentration with particles subject to probabilistic creation and annihilation. However, this method also suffers from the accumulation of per-step complexity over time. To address this limitation, Wang et al. [SIPF1] developed a stochastic interacting particle-field (SIPF) method that computes the chemo-attractant concentration through spectral decomposition and maintains only the current-step particle positions, which ensures that the per-step complexity remains constant over time.
However, we have to point out that in the SIPF method [SIPF1], each particle of $\rho$ interacts with every Fourier mode of $c$. This requires the computation of the trigonometric sum $\frac{1}{P}\sum_{p=1}^{P}e^{-\frac{2\pi i\mathbf{q}}{L}\cdot\widehat{\mathbf{X}}_{p}}$ for all wavenumbers $\mathbf{q}$, where $\widehat{\mathbf{X}}_{p}$ for $1\leq p\leq P$ are the positions of the particles (see Algorithm 1). For 3D systems with $H$ Fourier modes per dimension, this results in an $O(PH^{3})$ computational complexity per timestep. Such a computational bottleneck severely restricts the applicability of SIPF for large-scale 3D KS systems, particularly in scenarios requiring high-resolution and long-time simulations.
In this paper, to reduce the complexity, we propose the SIPF-PIC method (see Algorithm 4) that integrates localized particle-grid operations with spectral field solvers accelerated by the fast Fourier transform (FFT). This hybrid method confines particle interactions to adjacent grid nodes while resolving global frequency-domain couplings via FFT, yielding a computational complexity of $O\left(P+H^{3}\log H\right)$ per timestep, as presented in Theorem 3.7. The bidirectional interaction mechanism operates in each time step through two stages: particle-to-grid projection (see Algorithm 2) transfers particle-derived data to the computational grid, while grid-to-particle interpolation (see Algorithm 3) distributes field solutions from the grid back to individual particles; details are provided in Section 3. The particle-to-grid projection in the SIPF-PIC method, which numerically realizes the trigonometric summation that is also known as the unequally spaced fast Fourier transform (USFFT) [USFFT], has been employed in Fourier-basis particle simulations [USFFTS]. The grid-to-particle interpolation is employed in classical particle-in-cell methods [alma99, TskhakayaD.2007TPM] and does not rely on spectral decompositions. Our algorithm reduces the particle-grid interaction complexity from $O\left(PH^{3}\right)$ to $O\left(P+H^{3}\log H\right)$, thus enabling the simulation of 3D KS systems with significantly increased particle numbers $(P>10^{6})$ and finer grids $(H=256)$, and resulting in improved numerical accuracy.
[…]
To complement the theoretical analysis above, we also present numerical experiments to verify the convergence rate, computational complexity, and consistency with respect to $H$. We verify the critical mass $8\pi$ and the evolution of the second moment in a standard two-dimensional parabolic-elliptic system (1), where the dimension $d=2$ and the parameters $(k,\epsilon,\mu,\chi)=(0,0,1,1)$. For 3D parabolic-parabolic KS systems with $k,\epsilon>0$, the critical mass lacks a precise characterization due to its dependence on the initial density distribution $\rho(\mathbf{x},0)$. We thus validate the accuracy of the SIPF-PIC method under radially symmetric initial data, where the reference solution can be obtained by reducing the system to 1D and running a finite difference method on a very fine mesh.
Moreover, the SIPF-PIC method successfully reproduces complex blow-up patterns beyond the classical single-point radial type blow-up. To the best of our knowledge, this work presents the first numerical approach capable of resolving complex 3D blow-up dynamics beyond the single-point collapse regime, marking a key advance in the numerical simulation of such singular phenomena. We first investigate a non-radially symmetric configuration consisting of four uniform density balls located at the vertices of a tetrahedron, and then accurately resolve the ring-shaped blow-up pattern proposed in a recent paper [ring_ref], where mass concentrates on an evolving ring $x_{1}^{2}+x_{2}^{2}=r(t)^{2}$ with a time-dependent radius $r(t)$. In contrast, the original SIPF method, with parameter scalings ($P=10^{4},H=24$), lacks sufficient resolution to capture the ring-shaped blow-up pattern. Specifically, the SIPF-PIC method successfully captures the line-density blow-up, which is characterized by mass concentration along evolving spatial structures rather than isolated points, and further captures its instability under symmetry-breaking perturbations [collapsing]. These complex collapse mechanisms deviate fundamentally from classical delta-type blow-up dynamics, and the SIPF-PIC method, by leveraging large particle numbers and high-resolution grids, reliably captures their evolutionary processes in simulations. These results validate the performance of the SIPF-PIC method in addressing challenging 3D blow-up problems that were previously intractable by using existing methods.
The rest of the paper is organized as follows. Section 2 presents theoretical finite-time blow-up results for the KS system (1), and introduces the SIPF method, a hybrid particle-field method derived through temporal discretization of the system. In Section 3, we formalize the interpolation operators governing bidirectional particle-grid coupling, including particle-to-grid discretization and grid-to-particle interpolation, followed by the detailed formulation of the SIPF-PIC method. Subsection 4 rigorously establishes a comprehensive error analysis framework that quantifies the approximation errors from various sources in the SIPF-PIC method. Section 5 validates the method through numerical experiments, including convergence tests, performance benchmarks, and the visualization of 3D blow-up patterns. Finally, a summary of our findings and future research directions is given in Section 6.
2 Preliminary results
In this section, we present preliminary analytical and numerical results in the study of parabolic-parabolic KS system.
2.1 Analytical results of KS system
Recall the parabolic-parabolic KS system:
(2) $\left\{\begin{aligned} &\rho_{t}=\nabla\cdot\left(\mu\nabla\rho-\chi\rho\nabla c\right),\\ &\epsilon c_{t}=\Delta c-k^{2}c+\rho,&&\mathbf{x}\in\Omega,\ t\in[0,T],\end{aligned}\right$.
where $\rho(\mathbf{x},t)$ denotes the particle density, $c(\mathbf{x},t)$ the chemoattractant concentration, with $\chi,\mu>0$ (chemotactic sensitivity and diffusivity) and $\epsilon,k>0$ (relaxation and degradation rates).
Finite-Time Blow-Up Behaviors
For the parabolic-elliptic case (2) with $k=\epsilon=0$ and $\mu=1$, the spatial dimension $d$ governs the scaling properties of blow-up dynamics. In two dimensions ($d=2$), the system is $L^{1}$-critical: the scale transformation $\rho_{\lambda}(\mathbf{x},t)=\frac{1}{\lambda^{2}}\rho\left(\frac{\mathbf{x}}{\lambda},\frac{t}{\lambda^{2}}\right)$ maintains solution to system (2), and preserves the total mass $\int_{\mathbb{R}^{2}}\rho(\mathbf{x},t)\mathrm{d}\mathbf{x}$. There is a critical mass threshold $M_{\text{crit}}=\frac{8\pi}{\chi}$, independent of the initial density configuration. Subcritical initial distributions satisfying $\int_{\mathbb{R}^{2}}\rho(\mathbf{x},0)\,\mathrm{d}\mathbf{x}<M_{\text{crit}}$ admit global-in-time solutions [dolbeault2004optimal], whereas supercritical initial masses with $\int_{\mathbb{R}^{2}}\rho(\mathbf{x},0)\,\mathrm{d}\mathbf{x}>M_{\text{crit}}$ exhibit finite-time blow-up. In dimensions $d\geq 3$, the system becomes mass-supercritical, since the scaling $\rho_{\lambda}(\mathbf{x},t)=\frac{1}{\lambda^{d}}\rho\left(\frac{\mathbf{x}}{\lambda},\frac{t}{\lambda^{2}}\right)$ preserves the $L^{d/2}$ norm of $\rho$ instead of $L^{1}$. If the $L^{d/2}$ norm of the initial density configuration is small enough, i.e., $\|\rho(\mathbf{x},0)\|_{L^{d/2}}\leq N(d,\chi)$ for some dimension-dependent $N(d,\chi)$, the system admits a global-in-time solution with a decreasing $L^{d/2}$ norm [corrias2004global], that is, $\|\rho(\mathbf{x},t)\|_{L^{d/2}}\leq\|\rho(\mathbf{x},0)\|_{L^{d/2}}$ for all $t\geq 0$. Conversely, finite-time blow-up may occur [collapsing] for an arbitrary initial mass $M_{0}=\|\rho(\mathbf{x},0)\|_{L^{1}}$.
The blow-up is classified as Type I if
$\limsup_{t\to T^{-}}(T-t)\|\rho(t)\|_{L^{\infty}(\mathbb{R}^{d})}<C$
for some constant $C>0$. Otherwise, the blow-up is classified as Type II. Type II blow-up requires $d\geq 3$ and corresponds to singularities that do not collapse into a single-point delta measure, instead forming spatially distributed concentration patterns such as ring-shaped structures, as shown in [ring_ref].
Critical Mass
For the 2D parabolic-elliptic case in (2) with parameters $(k,\epsilon,\mu,\\ \chi)=(0,0,1,1)$, the critical mass threshold $M_{\text{crit}}=8\pi$ governs the formation of finite-time singularities. Solutions with initial mass $M_{0}>8\pi$ collapse into singular measures within finite time $T^{*}<\infty$; while subcritical masses $M_{0}<8\pi$ exhibit dispersion. The transition is characterized by the evolution of the second moment [Transport]:
(3) $\frac{\mathrm{d}}{\mathrm{d}t}m_{2}(t)=4M_{0}\left(1-\frac{M_{0}}{8\pi}\right),$
where
(4) $\quad m_{2}(t):=\int_{\mathbb{R}^{2}}|\mathbf{x}|^{2}\rho(t,\mathbf{x})\mathrm{d}\mathbf{x}.$
In other words, $m_{2}(t)$ grows linearly for $M_{0}<8\pi$, and decays linearly for $M_{0}>8\pi$. For $M_{0}>8\pi$, we conclude that $m_{2}(t)$ should become negative in finite time, which is impossible since $\rho$ is nonnegative. Therefore, a blow-up occurs before this contradiction arises. For further details on the critical mass analysis in the 2D parabolic-elliptic case, see [collapsing].
Unlike the simple 2D parabolic-elliptic KS system, where the total mass is the only factor that determines the aggregation behaviors, there are no similar results in more complex cases [he2022fast]. Numerical experiments [SIPF1] demonstrate that in 3D parabolic-parabolic KS systems, the critical mass depends on the initial distribution of $\rho$.
2.2 SIPF Method
In this section, we briefly introduce the Stochastic Interacting Particle Field (SIPF) method developed in [SIPF1]. The SIPF method is a hybrid particle-spectral approach, which approximates the solutions to (2) through coupled particle dynamics and field equations, avoiding historical dependence in drift computations while maintaining bounded operational complexity.
Throughout this section, the system is restricted to a domain $\Omega=\left[-\frac{L}{2},\frac{L}{2}\right]^{3}$ with periodic boundary conditions. In the SIPF method for solving the parabolic-parabolic KS system (2), the chemical concentration $\hat{c}(\mathbf{x},n\tau)$ is represented in frequency space using a Fourier series expansion, while the particle density $\hat{\rho}(\mathbf{x},n\tau)$ is represented in physical space by some moving particles. The temporal domain $[0,T]$ is partitioned into $N_{T}=\left\lceil\frac{T}{\tau}\right\rceil$ equal intervals of size $\tau$, where the ceiling function $\lceil x\rceil$ denotes the smallest integer greater than or equal to $x$.
At each time step $t_{n}=n\tau$, the concentration is represented as:
(5) $\hat{c}^{(n)}(\mathbf{x}):=\hat{c}(\mathbf{x},n\tau)=\sum_{\mathbf{q}\in U}\hat{\alpha}_{\mathbf{q}}^{(n)}e^{i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}},$
where $\hat{\alpha}_{\mathbf{q}}^{(n)}$ are the time-dependent Fourier coefficients for $\hat{c}$, and the frequency index set $U:=\left\{-\frac{H}{2},\ldots,\frac{H}{2}\right\}^{3}$. The particle density is represented as:
(6) $\hat{\rho}^{(n)}(\mathbf{x})=\frac{M_{0}}{P}\sum_{p=1}^{P}\delta\left(\mathbf{x}-\widehat{\mathbf{X}}_{p}^{(n)}\right),$
where $\delta(\cdot)$ is the Dirac delta function, $M_{0}=\int_{\Omega}\rho^{(0)}(\mathbf{x})\,\mathrm{d}\mathbf{x}$ is the initial mass, and $(\widehat{\mathbf{X}}_{p}^{(n)})_{1\leq p\leq P}$ are positions of particles at time $t_{n}$.
Updating chemical concentration $\hat{c}$
We discretize the governing equation for $c$ in (2) using an implicit Euler scheme:
(7) $\frac{\epsilon(\hat{c}^{(n)}-\hat{c}^{(n-1)})}{\tau}=\Delta\hat{c}-k^{2}\hat{c}+\hat{\rho}$.
Applying the Fourier transform to (7) yields:
(8) $\epsilon\frac{\hat{\alpha}_{\mathbf{q}}^{(n)}-\hat{\alpha}_{\mathbf{q}}^{(n-1)}}{\tau}=-\left(\frac{4\pi^{2}}{L^{2}}|\mathbf{q}|^{2}+k^{2}\right)\hat{\alpha}_{\mathbf{q}}^{(n)}+\hat{\beta}_{\mathbf{q}}^{(n-1)},$
where the Fourier coefficients for the density $\hat{\rho}$ at time $t=(n-1)\tau$, denoted by $\hat{\beta}_{\mathbf{q}}^{(n-1)}$, are computed as:
(9) $\hat{\beta}_{\mathbf{q}}^{(n-1)}=\frac{1}{L^{3}}\int_{\Omega}\hat{\rho}^{(n-1)}(\mathbf{x})e^{-i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}\,\mathrm{d}\mathbf{x}=\frac{M_{0}}{PL^{3}}\sum_{p=1}^{P}e^{-\frac{2\pi i}{L}\mathbf{q}\cdot\widehat{\mathbf{X}}_{p}^{(n-1)}}.$
The equation (8) simplifies to the update rule of $\hat{\alpha}_{\mathbf{q}}^{(n)}$:
(10) $\hat{\alpha}_{\mathbf{q}}^{(n)}=\frac{1}{1+\frac{\tau}{\epsilon}\left(\frac{4\pi^{2}}{L^{2}}|\mathbf{q}|^{2}+k^{2}\right)}\hat{\alpha}_{\mathbf{q}}^{(n-1)}+\frac{1}{\frac{4\pi^{2}}{L^{2}}|\mathbf{q}|^{2}+k^{2}+\frac{\epsilon}{\tau}}\hat{\beta}_{\mathbf{q}}^{(n-1)}.$
Updating particle density $\rho$
The governing equation for $\rho$ in (2) is the Fokker-Planck equation of the following SDE, where $\mathbf{W}_{t}=\sqrt{2\mu}\mathbf{B}_{t}$ and $\mathbf{B}_{t}$ is a standard Brownian motion:
(11) $\mathrm{d}\mathbf{X}=\chi\nabla_{\mathbf{X}}c(\mathbf{X},t)\mathrm{d}t+\mathrm{d}\mathbf{W}_{t}.$
We apply the Euler-Maruyama scheme:
(12) $\widehat{\mathbf{X}}_{p}^{(n)}=\widehat{\mathbf{X}}_{p}^{(n-1)}+\chi\tau\nabla\hat{c}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})+\mathbf{W}_{p}^{(n)},$
where $\mathbf{W}_{p}^{(n)}\sim\mathcal{N}(0,\,2\mu\tau I_{3})$ are independent 3D Wiener increments, $I_{3}$ is the $3\times 3$ identity matrix, and the drift $\nabla\hat{c}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})$ is computed by
(13) $\nabla\hat{c}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})=\sum_{\mathbf{q}\in U}\mathbf{q}\frac{2\pi i}{L}\hat{\alpha}^{(n-1)}_{\mathbf{q}}e^{\frac{2\pi i}{L}\mathbf{q}\cdot\widehat{\mathbf{X}}_{p}^{(n-1)}}.$
We summarize the SIPF method as Algorithm 1. We can find that the computational complexity per timestep is $O\left(PH^{3}\right)$, where operations (9) and (13) are the bottlenecks.
Algorithm 1 The SIPF method developed in [SIPF1]
1:Initialization:
[…]
3:Sample $\widehat{\mathbf{X}}_{p}^{(0)}\sim\text{i.i.d. }\mathbf{X}^{(0)}$. $\triangleright$ With density $\frac{\rho}{M_{0}}$ at $t=0$
4:
5:Update:
6:for $n=1,2,\dots,N_{T}$ do
7:  Update $\hat{\alpha}^{(n)}_{H\times H\times H}$:
8:  $\hat{\alpha}_{\mathbf{q}}^{(n)}\leftarrow\frac{1}{1+\frac{\tau}{\epsilon}\left(\frac{4\pi^{2}}{L^{2}}|\mathbf{q}|^{2}+k^{2}\right)}\hat{\alpha}_{\mathbf{q}}^{(n-1)}+\frac{1}{\frac{4\pi^{2}}{L^{2}}|\mathbf{q}|^{2}+k^{2}+\frac{\epsilon}{\tau}}\hat{\beta}_{\mathbf{q}}^{(n-1)}$,
9:  where $\hat{\beta}_{\mathbf{q}}^{(n-1)}=\frac{M_{0}}{PL^{3}}\sum_{p=1}^{P}e^{-\frac{2\pi i}{L}\mathbf{q}\cdot\widehat{\mathbf{X}}_{p}^{(n-1)}}$.
10:
11:  Update $\widehat{\mathbf{X}}_{p}^{(n)}$:
12:  for $p=1$ to $P$ do
13:   Compute $\mathbf{Z}_{p}^{(n)}\leftarrow\widehat{\mathbf{X}}_{p}^{(n-1)}+\chi\tau\nabla\hat{c}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})$.
14:   Sample $\mathbf{W}_{p}^{(n)}\sim\mathcal{N}(0,2\mu\tau I_{3})$. $\triangleright$ Independent noise
15:   Update $\widehat{\mathbf{X}}_{p}^{(n)}\leftarrow\mathbf{Z}_{p}^{(n)}+\mathbf{W}_{p}^{(n)}$.
16:  end for
17:      where $\nabla\hat{c}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})=\sum_{\mathbf{q}\in U}\mathbf{q}\frac{2\pi i}{L}\hat{\alpha}^{(n-1)}_{\mathbf{q}}e^{\frac{2\pi i}{L}\mathbf{q}\cdot\widehat{\mathbf{X}}_{p}^{(n-1)}}$.
18:end for
3 SIPF-PIC Method
Building on this computational bottleneck, we now introduce particle-in-cell operators as a pathway to a more scalable algorithm. The core idea is to leverage a structured Eulerian representation to reduce the complexity of particle-frequency interactions in the SIPF method, while retaining the Lagrangian flexibility of particles for transport and sampling.
3.1 Particle-in-cell interpolation framework
The drawback of the original SIPF method (Algorithm 1) is the particle-frequency interaction with $O(PH^{3})$ computational complexity. Our new framework is designed to break through this computational bottleneck.
The efficient interaction between the physical domain and the frequency domain is based on the FFT. We interpolate the particles using only a few nearby grid points, while the grid as a whole interacts with the frequency domain. This approach avoids the high complexity of interactions between each particle and each frequency.
In FFT, we require physical domain and frequency domain grids of the same size. We define the physical grid as $U\times\frac{L}{H}=\{-\frac{L}{2},\ldots,\frac{L}{2}-\frac{L}{H}\}^{3}$, where the index set $U$ is described in (5).
Consider a smooth target function $f:\mathbb{R}^{3}\to\mathbb{R}$; $\mathbf{X}\in\Omega$ is an arbitrary coordinate. A conventional particle-in-cell (PIC) method [alma99] represents $f(\mathbf{X})$ as a combination of the function values at grid points:
(14) $f(\mathbf{X})\approx\sum_{\mathbf{q}\in U}R\left(\mathbf{X},\mathbf{q}\frac{L}{H}\right)f\left(\mathbf{q}\frac{L}{H}\right),$
where the particle-grid interaction kernel $R\left(\mathbf{X},\mathbf{q}\frac{L}{H}\right)$ satisfies $\sum_{\mathbf{q}\in U}R\left(\mathbf{X},\mathbf{q}\frac{L}{H}\right)=1$ for every $\mathbf{X}\in\Omega$. Furthermore, the kernel $R(\mathbf{X},\cdot)$ is locally supported, meaning that for a given $\mathbf{X}\in\Omega$, only a few nearby grid points contribute to the approximation.
For example, in a trilinear interpolation with a second-order truncation error, the kernel is represented as
(15) $R\left(\mathbf{X},\mathbf{q}\frac{L}{H}\right)=\prod_{1\leq j\leq 3}\max\left(1-\frac{H}{L}\left|X_{j}-q_{j}\frac{L}{H}\right|,0\right).$
In particular, $R(\mathbf{X},\cdot)$ is nonzero only for the eight vertices of the grid cell containing $\mathbf{X}$. We denote the second-order basis $\Gamma_{2}=\{0,1\}^{3}$ and the relative grid positions $\mathbf{X}_{{\gamma}}=\left(\left\lfloor\frac{H}{L}\mathbf{X}\right\rfloor+{\gamma}\right)\times\frac{L}{H}$, where the notation $\lfloor\cdot\rfloor$ denotes the floor function applied component-wise. The interpolation (14) can be compactly expressed as
(16) $f(\mathbf{X})\approx\sum_{\boldsymbol{\gamma}\in\Gamma_{2}}F_{{\gamma}}^{(2)}(\boldsymbol{\lambda})f(\mathbf{X}_{{\gamma}}),$
where the weights $F_{j_{1}j_{2}j_{3}}^{(2)}(\boldsymbol{\lambda})=\prod_{s=1}^{3}\left(1-j_{s}+(2j_{s}-1)\lambda_{s}\right)$ for $(j_{1},j_{2},j_{3})\in\Gamma_{2}$, and $\boldsymbol{\lambda}=\boldsymbol{\lambda}(\mathbf{X}):=\frac{H}{L}\mathbf{X}-\left\lfloor\frac{H}{L}\mathbf{X}\right\rfloor\in[0,1)^{3}$ is the relative location of $\mathbf{X}$ in the grid cell containing $\mathbf{X}$. For points near the boundary under periodic boundary conditions, the indices are wrapped around the grid edges.
To achieve higher precision, we also introduce the fourth-order basis
$\Gamma_{4}=\Gamma_{2}\cup\{-1,2\}\times\{0,1\}\times\{0,1\}\cup\{0,1\}\times\{-1,2\}\times\{0,1\}\cup\{0,1\}\times\{0,1\}\times\{-1,2\},$
which enables interpolation [tricubic] exhibiting a fourth-order truncation error, as shown in Figure 1.
Figure 1: Extended vertices for fourth-order.
The complete fourth-order interpolation formula becomes as follows:
(17) $f(\mathbf{X})\approx\sum_{\gamma\in\Gamma_{4}}F_{\gamma}^{(4)}(\boldsymbol{\lambda})f(\mathbf{X}_{\gamma}),$
where the weights $F_{\gamma}^{(4)}(\boldsymbol{\lambda})$ are shown below, exhibiting cubic symmetry:
(18) $F_{2,j_{2},j_{3}}^{(4)}(\boldsymbol{\lambda})=-\frac{(1+\lambda_{1})(1-\lambda_{1})\lambda_{1}\left(1-j_{2}+(2j_{2}-1)\lambda_{2}\right)\left(1-j_{3}+(2j_{3}-1)\lambda_{3}\right)}{6},$
$F_{-1,j_{2},j_{3}}^{(4)}(\boldsymbol{\lambda})=-\frac{(2-\lambda_{1})\lambda_{1}(1-\lambda_{1})\left(1-j_{2}+(2j_{2}-1)\lambda_{2}\right)\left(1-j_{3}+(2j_{3}-1)\lambda_{3}\right)}{6},$
$\cdots$
$F_{j_{1}j_{2}j_{3}}^{(4)}(\boldsymbol{\lambda})=\prod_{s=1}^{3}\left(1-j_{s}+(2j_{s}-1)\lambda_{s}\right)\left(1+\sum_{k=1}^{3}\frac{\lambda_{k}(1-\lambda_{k})}{2}\right),(j_{1},j_{2},j_{3})\in\{0,1\}^{3}.$
We define particle-to-grid discretization as follows.
Definition 3.1.
[…]
(19) $I(\sum_{p=1}^{P}w_{p}\delta(\mathbf{x}-\widehat{\mathbf{X}}_{p}))=\sum_{p=1}^{P}w_{p}\sum_{\gamma\in\Gamma}F_{p,\gamma}\delta(\mathbf{x}-\widehat{\mathbf{X}}_{p,\gamma}),$
where $\Gamma$ is the interpolation basis index set from Section 3.1, $F_{p,\gamma}:=F_{\gamma}(\boldsymbol{\lambda}(\mathbf{X}_{p}))$ and $\widehat{\mathbf{X}}_{p,\gamma}:=\left(\left\lfloor\frac{H}{L}\mathbf{X}_{p}\right\rfloor+{\gamma}\right)\times\frac{L}{H}$ are the corresponding coefficients and grid points in (16) or (17) for $\widehat{\mathbf{X}}_{p}$.
Similarly, we define grid-to-particle interpolation as follows.
Definition 3.2.
The interpolation operator $I^{*}:\mathbb{R}^{H\times H\times H}\to C(\Omega)$ constructs continuous approximations via:
(20) $(I^{*}f)(\mathbf{X})=\sum_{\gamma\in\Gamma}F_{\gamma}(\boldsymbol{\lambda}(\mathbf{X}))f(\mathbf{X}_{\gamma}),$
where $f$ is sampled at grid points $\mathbf{X}_{\gamma}$.
The operator $I^{*}$ produces piecewise polynomial functions that depend only on the nodal values, satisfying the adjoint relation $\langle I\rho,f\rangle=\langle\rho,I^{*}f\rangle$ for $\rho\in\mathcal{M}(\Omega)$ and $f$ bounded.
For completeness, we consider the following convention of the discrete Fourier transform $\mathcal{F}_{H}$, which admits the following definition:
(21) $\mathcal{F}_{H}f\left(\mathbf{y}\right):=L^{-3}\int_{\mathbf{x}\in\Omega}e^{-i\mathbf{x}\cdot\mathbf{y}}f(\mathbf{x})\,\mathrm{d}\mathbf{x},$
where $\mathbf{y}=\mathbf{q}\frac{2\pi}{L}$ is a frequency vector. The inverse Fourier transform operator $\mathcal{F}_{H}^{-1}$ is then given by:
(22) $\mathcal{F}_{H}^{-1}\hat{\alpha}(\mathbf{x})=\sum_{\mathbf{q}\in U}\hat{\alpha}_{\mathbf{q}}e^{i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}.$
The Fourier coefficients $\hat{\alpha}_{\mathbf{q}}$ and $\hat{\beta}_{\mathbf{q}}$ defined in (5) and (9) are specific evaluations: $\hat{\alpha}_{\mathbf{q}}=\mathcal{F}_{H}\hat{c}\left(\mathbf{q}\frac{2\pi}{L}\right),\quad\hat{\beta}_{\mathbf{q}}=\mathcal{F}_{H}\hat{\rho}\left(\mathbf{q}\frac{2\pi}{L}\right).$ The composition $\mathcal{F}_{H}^{-1}\circ\mathcal{F}_{H}$ forms a projection onto the space of trigonometric polynomials $\mathcal{T}_{H}:=\left\{f:f=\sum_{\mathbf{q}\in U}\hat{\alpha}_{\mathbf{q}}e^{i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}\right\}$. For any function $f\in\mathcal{T}_{H}$, it holds that $\mathcal{F}_{H}^{-1}\circ\mathcal{F}_{H}(f)=f.$ Additionally, for $\mathbf{q},\mathbf{q^{\prime}}\in U$ with $\mathbf{q}\neq\mathbf{q^{\prime}}$, the following properties are satisfied:
(23) $\mathcal{F}_{H}^{-1}\circ\mathcal{F}_{H}\left(\delta\left(\mathbf{x}-\mathbf{q}\frac{L}{H}\right)\right)\left(\mathbf{q}\frac{L}{H}\right)$ $=\left(\frac{H}{L}\right)^{3},$
$\mathcal{F}_{H}^{-1}\circ\mathcal{F}_{H}\left(\delta\left(\mathbf{x}-\mathbf{q}\frac{L}{H}\right)\right)\left(\mathbf{q^{\prime}}\frac{L}{H}\right)$ $=0.$
3.2 Computational complexity and algorithmic details
Consider $P$
equally weighted particles at positions $\{\widehat{\mathbf{X}}_{p}\}_{p=1}^{P}\subset\Omega$ with total mass $M_{0}$. The empirical measure is shown as:
(24) $\hat{\rho}(\mathbf{x})=\frac{M_{0}}{P}\sum_{p=1}^{P}\delta(\mathbf{x}-\widehat{\mathbf{X}}_{p}).$
We set the computational grid $\mathcal{L}=\{\mathbf{X}_{\gamma}=\frac{L}{H}\mathbf{q}:\mathbf{q}\in U\}$ with spacing $h=\frac{L}{H}$. We perform Algorithm 2 to compute $\mathcal{F}_{H}I(\hat{\rho})(\mathbf{y})$ for all $\mathbf{y}\in U\times\frac{2\pi}{L}$, where the discretization operator $I$ is specified in Definition 3.1.
Algorithm 2 particles-to-grid discretization
1:Initialize zero tensor $\phi[\mathbf{q}]:=\phi_{\mathbf{q}}$ for all $\mathbf{q}\in U$
2:for $p=1$ to $P$ do
3:  for $\gamma\in\Gamma$ do $\triangleright$ Defined in Section 3.1
4:   Find grid index: $\mathbf{q}_{p,\gamma}=\mathbf{X}_{p,\gamma}\cdot\frac{H}{L}$
5:   Calculate interpolation weight: $F_{p,\gamma}$
6:   Update density: $\phi[\mathbf{q}_{p,\gamma}]\leftarrow\phi[\mathbf{q}_{p,\gamma}]+\frac{M_{0}F_{p,\gamma}}{P}$
7:  end for
8:end for
9:Compute $\mathcal{F}_{H}\phi$ via Fast Fourier Transform
Lemma 3.3.
With $H$ Fourier modes and $P$ particles, the complexity of particle-to-grid discretization in Algorithm 2 is $O(H^{3}\log H+P)$.
Proof 3.4.
The particle-to-grid transfer costs $O(P)$ (each particle contributes to $|\Gamma|$ neighbors), and the FFT on the $H\times H\times H$ grid costs $O(H^{3}\log H)$.
For the inverse operation, given the values of a function $f$ on the grid $f\left(U\cdot\frac{L}{H}\right)$, we perform Algorithm 3 to compute $I^{*}f(\mathbf{X}_{1}),\ldots,I^{*}f(\mathbf{X}_{P})$, where the interpolation operator $I^{*}$ is specified in Definition 3.2.
Algorithm 3 grid-to-particles interpolation
1:Input tensor $f[\mathbf{q}]:=f_{\mathbf{q}}$ defined on grid indices $\mathbf{q}\in U$
2:for $p=1$ to $P$ do
3:  Initialize particle value: $f_{p}\leftarrow 0$
4:  for $\gamma\in\Gamma$ do $\triangleright$ Defined in Section 3.1
5:   Find grid index: $\mathbf{q}_{p,\gamma}=\mathbf{X}_{p,\gamma}\cdot\frac{H}{L}$
6:   Obtain interpolation weight: $F_{p,\gamma}$
7:   Accumulate contribution: $f_{p}\leftarrow f_{p}+F_{p,\gamma}\cdot f[\mathbf{q}_{p,\gamma}]$
8:  end for
9:  Store interpolated value: $I^{*}f(\mathbf{X}_{p})\leftarrow f_{p}$
10:end for
Lemma 3.5.
With $P$ particles, the complexity of the grid-to-particle interpolation in Algorithm 3 is $O(P)$.
Proof 3.6.
The grid-to-particle transfer costs $O(P)$ (each particle interacts with $|\Gamma|$ neighbors).
The computational complexity of the original method, dominated by all-to-all particle-frequency interactions scaling as $O(PH^{3})$, is reduced to $O(P+H^{3}\log H)$ via the PIC acceleration presented in Section 3.1. This acceleration involves local grid-projected density interpolation and gradient reconstruction, limiting particle interactions to neighboring grid points while handling frequency components globally. Periodic boundary conditions are imposed on the boundary grids. To enhance numerical stability, we implement a spectral cutoff that filters out high-frequency modes exceeding the threshold $H_{0}\cdot\frac{2\pi}{L}$.
Algorithm 4 SIPF-PIC method
1:Initialization:
[…]
3:Sample $\widehat{\mathbf{X}}_{p}^{(0)}\sim\text{i.i.d. }\mathbf{X}^{(0)}$
4:
5:Update:
6:for $n=1,2,\dots,N_{T}$ do
7:  Update $\hat{\alpha}_{\mathbf{q}}^{(n)}$:
8:  $\hat{\alpha}_{\mathbf{q}}^{(n)}\leftarrow\frac{1}{1+\frac{\tau}{\epsilon}\left(\frac{4\pi^{2}}{L^{2}}|\mathbf{q}|^{2}+k^{2}\right)}\hat{\alpha}_{\mathbf{q}}^{(n-1)}+\frac{1}{\frac{4\pi^{2}}{L^{2}}|\mathbf{q}|^{2}+k^{2}+\frac{\epsilon}{\tau}}\mathcal{F}_{H}I(\hat{\rho}^{(n-1)})\left(\mathbf{q}\frac{2\pi}{L}\right)$
9:  where $\hat{\rho}^{(n-1)}(\mathbf{x})=\sum_{p=1}^{P}\frac{M_{0}}{P}\delta(\mathbf{x}-\widehat{\mathbf{X}}_{p}^{(n-1)})$
10:  and $\mathcal{F}_{H}I(\hat{\rho}^{(n-1)})\left(\mathbf{q}\frac{2\pi}{L}\right)=\frac{M_{0}}{PL^{3}}\sum_{p=1}^{P}e^{-\frac{2\pi i}{L}\mathbf{q}\cdot\widehat{\mathbf{X}}_{p}^{(n-1)}}$ computed by Alg. 2
11:
12:  Update $\widehat{\mathbf{X}}_{p}^{(n)}$:
13:  for $p=1$ to $P$ do
14:   Compute $\mathbf{Z}_{p}^{(n)}\leftarrow\widehat{\mathbf{X}}_{p}^{(n-1)}+\chi\tau I^{*}(\nabla\hat{c}^{(n-1)})(\widehat{\mathbf{X}}_{p}^{(n-1)})$
15:   Sample $\mathbf{W}_{p}^{(n)}\sim\mathcal{N}(0,2\mu\tau I_{3})$ $\triangleright$ 3D Brownian motion
16:   Update $\widehat{\mathbf{X}}_{p}^{(n)}\leftarrow\mathbf{Z}_{p}^{(n)}+\mathbf{W}_{p}^{(n)}$ $\triangleright$ Periodic setting
17:  end for
18:  with $\nabla\hat{c}^{(n-1)}\left(\mathbf{p}\frac{L}{H}\right)=\mathcal{F}_{H}^{-1}\left(\frac{2\pi i}{L}\mathbf{q}\hat{\alpha}_{\mathbf{q}}^{(n-1)}\mathbf{1}_{\{\|\mathbf{q}\|\leq H_{0}\}}\right)\left(\mathbf{p}\frac{L}{H}\right)$
[…]
20:  and $I^{*}(\nabla\hat{c}^{(n-1)})$ computed by Alg. 3
21:end for
Theorem 3.7.
The single-step complexity of Algorithm 4 is $O\left(H^{3}\log H+P\right)$.
Proof 3.8.
Each iteration costs $O(H^{3}\log H+P)$: spectral updates require FFT with complexity $O(H^{3}\log H)$ and particle-to-grid transfer with $O(P)$, while particle updates require inverse FFT with $O(H^{3}\log H)$ and gradient operations with $O(P)$.
3.3 Bounds of interpolation error
Recall the empirical measure (24) and the discretization operator $I$ defined in Definition 3.1. We claim the following error bounds for $I$ under a Fourier transform.
Lemma 3.9.
For any particle distribution $\{\mathbf{X}_{p}\}_{p=1}^{P}\subset\Omega$, the empirical density $\hat{\rho}^{(n-1)}$ is defined by (6). let the frequencies be $\mathbf{y}=\mathbf{q}\frac{2\pi}{L}$ for indices $\mathbf{q}\in U$. The interpolation errors satisfy:
(25) $|(\mathcal{F}_{H}I_{2}(\hat{\rho})-\mathcal{F}_{H}\hat{\rho})(\mathbf{y})|$ $\leq M_{0}C_{2}L^{-3}|\mathbf{y}|^{2}\left(\frac{L}{H}\right)^{2},$
(26) $|(\mathcal{F}_{H}I_{4}(\hat{\rho})-\mathcal{F}_{H}\hat{\rho})(\mathbf{y})|$ $\leq M_{0}C_{4}L^{-3}|\mathbf{y}|^{4}\left(\frac{L}{H}\right)^{4},$
where $C_{\eta}$ depends on the interpolation order $\eta$ with $\eta=2,4$.
Proof 3.10.
Let $g_{\mathbf{y}}(\mathbf{x})=e^{-i\mathbf{y}\cdot\mathbf{x}}$. Then we have the Fourier transform
(27) $\mathcal{F}_{H}\hat{\rho}(\mathbf{y})=\frac{M_{0}}{PL^{3}}\sum_{p=1}^{P}g_{\mathbf{y}}(\mathbf{X}_{p}),\quad\mathcal{F}_{H}I(\hat{\rho})(\mathbf{y})=\frac{M_{0}}{PL^{3}}\sum_{p=1}^{P}\sum_{\gamma\in\Gamma}F_{p,\gamma}g_{\mathbf{y}}(\mathbf{X}_{p,\gamma}).$
Using a $d$-th order Taylor expansion with Lagrange remainder at $\mathbf{X}$, for multi-index $\mathbf{k}=(k_{1},k_{2},k_{3})$ and $D_{\mathbf{k}}=\partial^{|\mathbf{k}|}/\partial x_{1}^{k_{1}}\partial x_{2}^{k_{2}}\partial x_{3}^{k_{3}}$, we have
(28) $g_{\mathbf{y}}(\mathbf{x})=\sum_{|\mathbf{k}|\leq\eta-1}\frac{D_{\mathbf{k}}g_{\mathbf{y}}(\mathbf{X})}{\mathbf{k}!}(\mathbf{x}-\mathbf{X})^{\mathbf{k}}+\sum_{|\mathbf{k}|=\eta}\frac{D_{\mathbf{k}}g_{\mathbf{y}}(\mathbf{X}+\xi)}{\mathbf{k}!}(\mathbf{x}-\mathbf{X})^{\mathbf{k}},$
where $\xi$ lies between $\mathbf{X}$ and $\mathbf{x}$, and $\mathbf{k}!=k_{1}!k_{2}!k_{3}!$. Throughout this paper, the notation $|\mathbf{k}|=k_{1}+k_{2}+k_{3}$ means the order of a multi-index $\mathbf{k}$, which is different from the norm $|\mathbf{x}|$ for a vector $\mathbf{x}$ in Euclidean space.
The exactness of $I_{\eta}$ for polynomials of degrees 0 to $\eta-1$ implies cancellation of lower-order terms, in other words, $\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}f(\mathbf{X}_{\gamma})=f(\mathbf{X})$ holds for all polynomials $f$ of degree at most $\eta-1$, where $\overline{F_{\gamma}}:=F_{\gamma}(\boldsymbol{\lambda}(\mathbf{X}))$, yielding:
(29) $g_{\mathbf{y}}(\mathbf{X})+\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}g_{\mathbf{y}}(\mathbf{X}_{\gamma})=\sum_{|\mathbf{k}|=\eta}\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}\frac{D_{\mathbf{k}}g_{\mathbf{y}}(\mathbf{X}+\xi(\mathbf{X},\gamma))}{\mathbf{k}!}(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}.$
Since $g_{\mathbf{y}}(\mathbf{x})=e^{-i\mathbf{y}\cdot\mathbf{x}}$, we have $|D_{\mathbf{k}}g_{\mathbf{y}}(\mathbf{X}+\xi(\mathbf{X},\gamma))|\leq|\mathbf{y}|^{\eta},$ for all $\mathbf{k},\mathbf{X},\xi,\gamma$.
For the case of $\eta=2$ and $\Gamma=\Gamma_{2}$, we have $\overline{F_{\gamma}}\geq 0$, $\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}=1$ and $(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}\leq\left(\frac{L}{H}\right)^{2}.$ Thus, we establish the second-order error bound
(30) $\left|-g_{\mathbf{y}}(\mathbf{X})+\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}g_{\mathbf{y}}(\mathbf{X}_{\gamma})\right|\leq\sum_{|\mathbf{k}|=2}\frac{|\mathbf{y}|^{2}}{\mathbf{k}!}\left(\frac{L}{H}\right)^{2}=\frac{9}{2}|\mathbf{y}|^{2}\left(\frac{L}{H}\right)^{2}.$
For the case of $\eta=4$, let $\Gamma=\Gamma_{4}$. The coefficients in the outer set $\Gamma\setminus\Gamma_{2}$ are negative, while the coefficients in the inner set $\Gamma_{2}$ are positive.
We only need to handle the following
(31) $\left|\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}\right|$ $\leq\left|\sum_{\gamma\in\Gamma_{2}}\overline{F_{\gamma}}(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}\right|+\left|\sum_{\gamma\in\Gamma\setminus\Gamma_{2}}\overline{F_{\gamma}}(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}\right|$
$\leq\left(\frac{L}{H}\right)^{4}\left(\sum_{\gamma\in\Gamma_{2}}\overline{F_{\gamma}}-2^{4}\sum_{\gamma\in\Gamma\setminus\Gamma_{2}}\overline{F_{\gamma}}\right)$
$=\left(\frac{L}{H}\right)^{4}\left(1-17\sum_{\gamma\in\Gamma\setminus\Gamma_{2}}\overline{F_{\gamma}}\right).$
We partition the 24 outer grid points into 8 clusters based on their nearest inner vertices. For instance, the cluster containing $\mathbf{X}_{211},\mathbf{X}_{121},$ and $\mathbf{X}_{112}$ shares adjacency with the inner vertex $\mathbf{X}_{111}$. For each cluster, the coefficient combination satisfies
(32) $(\overline{F}_{211}+\overline{F}_{121}+\overline{F}_{112})=\lambda_{1}\lambda_{2}\lambda_{3}\left(\frac{3-\lambda_{1}^{2}-\lambda_{2}^{2}-\lambda_{3}^{2}}{6}\right)\leq\frac{1}{2}\lambda_{1}\lambda_{2}\lambda_{3}.$
Aggregating over all 8 clusters yields
(33) $\sum_{\gamma\in\Gamma\setminus\Gamma_{2}}\overline{F_{\gamma}}\leq\frac{1}{2}\left(\lambda_{1}\lambda_{2}\lambda_{3}+\cdots+(1-\lambda_{1})(1-\lambda_{2})(1-\lambda_{3})\right)=\frac{1}{2}.$
Combining these results, we establish the fourth-order error bound
(34) $\left|-g_{\mathbf{y}}(\mathbf{X})+\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}g_{\mathbf{y}}(\mathbf{X}_{\gamma})\right|\leq\frac{19}{2}\sum_{|\mathbf{k}|=4}\frac{|\mathbf{y}|^{4}}{\mathbf{k}!}\left(\frac{L}{H}\right)^{4}=\frac{513}{16}|\mathbf{y}|^{4}\left(\frac{L}{H}\right)^{4}.$
The overall spectral error consequently satisfies
(35) $\left|\mathcal{F}_{H}I(\hat{\rho})(\mathbf{y})-\mathcal{F}_{H}\hat{\rho}(\mathbf{y})\right|$ $=\left|\frac{M_{0}}{PL^{3}}\left(\sum_{p=1}^{P}\sum_{\gamma\in\Gamma}F_{p,\gamma}g_{\mathbf{y}}(\mathbf{X}_{p,\gamma})-\sum_{p=1}^{P}g_{\mathbf{y}}(\mathbf{X}_{p})\right)\right|$
$\leq\frac{M_{0}}{PL^{3}}\cdot P\cdot C_{\eta}|\mathbf{y}|^{\eta}\left(\frac{L}{H}\right)^{\eta}$
$=C_{\eta}M_{0}L^{-3}|\mathbf{y}|^{\eta}\left(\frac{L}{H}\right)^{\eta},$
where $\eta=2$ (with $C_{2}=\frac{9}{2}$) and $\eta=4$ (with $C_{4}=\frac{513}{16}$) correspond to second-order and fourth-order interpolation, respectively.
As a direct consequence of Lemma 3.9 (see equations (30) and (34)), we present the error bounds for the operation $I^{*}$ specified in Definition 3.2 as follows.
Lemma 3.11 (Error bounds of grid-to-particle interpolation).
Assume $f\in\\ C^{2}(\Omega)$. Given the values of $f$ on the grid $f\left(U\cdot\frac{L}{H}\right)$, for each p=1,,Pp=1,...,P, the error of the second-order interpolation scheme at position $\mathbf{X}_{p}$ satisfies
(36) $|I^{*}_{2}f(\mathbf{X}_{p})-f(\mathbf{X}_{p})|\leq K_{2}\left(\frac{L}{H}\right)^{2};$
if we further assume $f\in C^{4}(\Omega)$, the error of the fourth-order interpolation scheme satisfies
(37) $|I^{*}_{4}f(\mathbf{X}_{p})-f(\mathbf{X}_{p})|\leq K_{4}\left(\frac{L}{H}\right)^{4},$
where $K_{2}=\frac{9}{2}\max_{|\mathbf{k}|=2}\|D_{\mathbf{k}}f\|_{\infty}$ is associated with the second derivative of $f$, and $K_{4}=\frac{513}{16}\max_{|\mathbf{k}|=4}\|D_{\mathbf{k}}f\|_{\infty}$ is associated with the fourth derivative of $f$.
4 Convergence of the SIPF-PIC method
In this section, we present the convergence result of the proposed SIPF-PIC method (Algorithm 4), which is stated as Theorem 4.16. Specifically, we introduce the following Brownian coupling technique to construct an auxiliary random variable $\widetilde{\mathbf{X}}$ that bridges the ground truth random variable $\mathbf{X}$ and the empirical random variable $\widehat{\mathbf{X}}$ maintained in the algorithm.
Definition 4.1.
The auxiliary particles $\widetilde{\mathbf{X}}_{p}^{(n)}$ (for $1\leq p\leq P$) satisfy the following update process:
(38) $\widetilde{\mathbf{X}}_{p}^{(0)}$ $=\mathbf{X}_{p}^{(0)},$
$\widetilde{\mathbf{X}}_{p}^{(n)}$ $=\widetilde{\mathbf{X}}_{p}^{(n-1)}+\chi\tau\nabla c(\widetilde{\mathbf{X}}_{p}^{(n-1)},(n-1)\tau)+\mathbf{W}_{p}^{(n)}.$
$\widetilde{\mathbf{X}}^{(n)}$ denotes the discrete random variable consisting of particles $\widetilde{\mathbf{X}}_{1}^{(n)},\ldots,\widetilde{\mathbf{X}}_{P}^{(n)}$, where each particle is sampled with probability $\frac{1}{P}$.
The auxiliary discrete density $\tilde{\rho}^{(n)}$ and its Fourier coefficients $\tilde{\beta}^{(n)}$ are defined as
(39) $\tilde{\rho}^{(n)}(\mathbf{x})=\frac{M_{0}}{P}\sum\limits_{p=1}^{P}\delta(\mathbf{x}-\widetilde{\mathbf{X}}_{p}^{(n)}),$
(40) $\tilde{\beta}_{\mathbf{q}}^{(n)}=\frac{M_{0}}{P}L^{-3}\sum_{p=1}^{P}e^{-i\widetilde{\mathbf{X}}_{p}^{(n)}\cdot\frac{2\pi}{L}\mathbf{q}}.$
Furthermore, the auxiliary concentration $\tilde{c}^{(n)}$ and its Fourier coefficients $\tilde{\alpha}^{(n)}$ are updated via Algorithm 4:
(41) $\tilde{\alpha}^{(n)}_{\mathbf{q}}=\frac{1}{1+\frac{\tau}{\epsilon}(\frac{4\pi^{2}}{L^{2}}|\mathbf{q}|^{2}+k^{2})}\tilde{\alpha}^{(n-1)}_{\mathbf{q}}+\frac{1}{\frac{4\pi^{2}}{L^{2}}|\mathbf{q}|^{2}+k^{2}+\frac{\epsilon}{\tau}}\tilde{\beta}^{(n-1)}_{\mathbf{q}}$,
(42) $\tilde{c}^{(n)}(\mathbf{x})=\sum_{\mathbf{q}\in U,\|\mathbf{q}\|\leq H_{0}}\tilde{\alpha}_{\mathbf{q}}^{(n)}e^{\frac{2\pi i}{L}\mathbf{q}\cdot\mathbf{x}}.$
The semi-discrete implicit reference concentration $c^{(n)}$ and its Fourier coefficients $\alpha^{(n)}$ are updated as follows:
(43) ${\alpha}^{(n)}_{\mathbf{q}}=\frac{1}{1+\frac{\tau}{\epsilon}(\frac{4\pi^{2}}{L^{2}}|\mathbf{q}|^{2}+k^{2})}{\alpha}^{(n-1)}_{\mathbf{q}}+\frac{1}{\frac{4\pi^{2}}{L^{2}}|\mathbf{q}|^{2}+k^{2}+\frac{\epsilon}{\tau}}{\beta}^{(n-1)}_{\mathbf{q}}$,
(44) ${\beta}_{\mathbf{q}}^{(n-1)}=\frac{1}{L^{3}}\int_{\Omega}{\rho}^{(n-1)}(\mathbf{x})e^{-i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}\,\mathrm{d}\mathbf{x},$
where the corresponding random variable $\mathbf{X}^{(n)}$ is updated via an Euler-Maruyama semi-discretization over the time interval, and $\rho^{(n)}$ denotes its distribution
(45) $\mathbf{X}^{(n)}=\mathbf{X}^{(n-1)}+\chi\tau\nabla{c}(\mathbf{X}^{(n-1)},(n-1)\tau)+\mathbf{W}^{(n)}.$
[…]
Remark 4.2.
The probability space in the algorithm is generated as follows: the initial values $(\mathbf{X}_{p}^{(0)})_{1\leq p\leq P}$, and the diffusion terms $(\mathbf{W}_{p}^{(n)})_{1\leq p\leq P}$ in the $n$-th step for $1\leq n\leq N_{T}$. When discussing probabilities or using the expectation notation $\mathbb{E}$, we always operate within this space. The updates of particles $\widehat{\mathbf{X}}_{p}^{(n)}$ and auxiliary particles $\widetilde{\mathbf{X}}_{p}^{(n)}$ depend on this shared probability space, which constitutes the essence of the Brownian coupling technique.
Definition 4.3.
The mean absolute error $d^{(n)}$ is defined as
(46) $d^{(n)}=\frac{1}{P}\sum\limits_{p=1}^{P}|\widehat{\mathbf{X}}_{p}^{(n)}-\widetilde{\mathbf{X}}_{p}^{(n)}|.$
The frequency error $e_{\mathbf{q}}^{(n)}$ is defined as $e_{\mathbf{q}}^{(n)}:=\hat{\alpha}_{\mathbf{q}}^{(n)}-\tilde{\alpha}_{\mathbf{q}}^{(n)}$, where the Fourier coefficients $\hat{\alpha}_{\mathbf{q}}$ and $\tilde{\alpha}_{\mathbf{q}}$ are defined via (10) and (41) for indices $\mathbf{q}\in U=\left\{-\frac{H}{2},\ldots,\frac{H}{2}-1\right\}^{3}$.
Remark 4.4.
The quantity $d^{(n)}$ serves as an upper bound for the $\mathcal{W}_{1}$ distance between the empirical distribution $\widehat{\mathbf{X}}^{(n)}$ and the auxiliary distribution $\widetilde{\mathbf{X}}^{(n)}$.
To ensure the convergence of Algorithm 4, we impose the following assumption.
Assumption 4.5.
The solution $\rho,c$ to (1) is periodic, that is, for $q_{1},q_{2},q_{3}\in\mathbf{Z}$,
(47) $\rho(x_{1}+q_{1}L,x_{2}+q_{2}L,x_{3}+q_{3}L)$ $=\rho(x_{1},x_{2},x_{3}),$
$c(x_{1}+q_{1}L,x_{2}+q_{2}L,x_{3}+q_{3}L)$ $=c(x_{1},x_{2},x_{3}).$
The function $\rho$ and its gradient $\nabla\rho$ are square-integrable, i.e.,
(48) $\int_{\mathbf{x}\in[-\frac{L}{2},\frac{L}{2}]^{3}}\rho(\mathbf{x},t)^{2}\,\mathrm{d}\mathbf{x}\leq N_{2}\quad\forall t\in\left[0,T\right],$
(49) $\int_{\mathbf{x}\in[-\frac{L}{2},\frac{L}{2}]^{3}}|\nabla\rho(\mathbf{x},t)|^{2}\,\mathrm{d}\mathbf{x}\leq N_{1,2}\quad\forall t\in\left[0,T\right].$
The first to third order spatial derivatives of $c$ and $\hat{c}$ are bounded, i.e., for any spatial multi-index $\mathbf{k}$ such that $|\mathbf{k}|=\eta$ where $\eta=1,2,3$, it holds that
(50) $\max(|D_{\mathbf{k}}c(\mathbf{x},t)|,|D_{\mathbf{k}}\hat{c}(\mathbf{x},t)|)\leq M_{\eta}\quad\forall\mathbf{x}\in\left[-\frac{L}{2},\frac{L}{2}\right]^{3},t\in\left[0,T\right].$
The temporal derivative of $\nabla c$ is bounded, that is,
(51) $\left|\frac{\partial\nabla c(\mathbf{x},t)}{\partial t}\right|\leq J_{1}\quad\forall\mathbf{x}\in\left[-\frac{L}{2},\frac{L}{2}\right]^{3},t\in\left[0,T\right].$
4.1 Main convergence results
We outline the structure of the argument. In this section, we first present a series of lemmas that establish bounds for individual components of the error dynamics, and then prove Theorem 4.16, which constitutes the main convergence result of Algorithm 4. Then, in Section 4.2, we further prove that under specific parameter combinations, the smoothness assumption on $\hat{c}$ in Assumption 4.5 can be removed.
The following lemma is the discrete Gronwall inequality.
Lemma 4.6.
Assume we have the growth estimate
(52) $d^{(n)}\leq d^{(n-1)}(1+\chi\tau M_{2})+\chi\tau F^{(n-1)}$
for some growth constant $M_{2}$ and one-step error $F^{(n)}$, and that we can achieve the one-step error control
(53) $F^{(n-1)}\leq G^{(n-1)}+K\max(d^{(0)},\ldots,d^{(n-1)})$
for some local truncation error $G^{(n)}$ and error amplification constant $K$. Then, the final mean absolute error $d^{\left(N_{T}\right)}$ satisfies
(54) $d^{\left(N_{T}\right)}\leq\left(1+(K+\chi M_{2})\tau\right)^{N_{T}}\left(d^{(0)}+\chi\tau\sum_{m=0}^{N_{T}-1}G^{(m)}\right).$
Proof 4.7.
We proceed by induction on $n$. For $n=0$, the bound holds trivially. Assume the inequality holds for $n=k$, that is,
(55) $d^{(k)}\leq\left(1+(K+\chi M_{2})\tau\right)^{k}\left(d^{(0)}+\chi\tau\sum_{m=0}^{k-1}G^{(m)}\right).$
From the growth estimate in the hypothesis,
(56) $d^{(k+1)}\leq d^{(k)}(1+\chi\tau M_{2})+\chi\tau F^{(k)}.$
By the induction hypothesis and the one-step error control $F^{(k)}\leq G^{(k)}+K\max(d^{(0)},\\ \ldots,d^{(k)})$, we first substitute these into the inequality for $d^{(k+1)}$ and then simplify to derive
(57) $d^{(k+1)}\leq\left(1+(K+\chi M_{2})\tau\right)^{k+1}\left(d^{(0)}+\chi\tau\sum_{m=0}^{k}G^{(m)}\right).$
Set $n=N_{T}$ to obtain the desired result.
We first estimate the single-step growth of $d^{(n)}$ as follows.
Lemma 4.8.
The mean absolute error $d^{(n)}$ is controlled by $d^{(n-1)}$ with a probability of at least $1-e^{-\frac{S_{1}^{2}}{2}}$:
(58) $d^{(n)}$ $\leq d^{(n-1)}(1+\chi\tau M_{2})+\chi\tau\Biggl(M_{3}\left(\frac{L}{H}\right)^{2}+2P^{-\frac{1}{2}}S_{1}M_{1}$
$\quad+\frac{N_{2}^{\frac{1}{2}}}{M_{0}}\left(\int_{\Omega}\left|\nabla c^{(n-1)}(\mathbf{x})-\nabla\hat{c}^{(n-1)}(\mathbf{x})\right|^{2}\mathrm{d}\mathbf{x}\right)^{\frac{1}{2}}\Biggr),$
where $M_{3}$ and $N_{2}$ are defined in Assumption 4.5, $S_{1}>0$ is a parameter controlling the probability bound, and $\nabla c^{(n-1)}=\nabla c(\cdot,(n-1)\tau)$.
Proof 4.9.
Define the mappings $g^{(n-1)}(\mathbf{X})=\mathbf{X}+\chi\tau\nabla c^{(n-1)}(\mathbf{X})$, $\overline{g}^{(n-1)}(\mathbf{X})=\mathbf{X}+\chi\tau\nabla\hat{c}^{(n-1)}(\mathbf{X})$, and $\hat{g}^{(n-1)}(\mathbf{X})=\mathbf{X}+\chi\tau I^{*}\nabla\hat{c}^{(n-1)}(\mathbf{X})$, respectively. Since the updates of the solutions $\mathbf{X}^{(n)}$ and $\widehat{\mathbf{X}}^{(n)}$ use the same diffusion $\mathbf{W}^{(n)}$, we have by the triangle inequality,
(59) $d^{(n)}$ $=\frac{1}{P}\sum_{p=1}^{P}\bigl|g^{(n-1)}(\widetilde{\mathbf{X}}_{p}^{(n-1)})-\hat{g}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})\bigr|$
$\leq\frac{1}{P}\left(\sum_{p=1}^{P}\bigl|g^{(n-1)}(\widetilde{\mathbf{X}}_{p}^{(n-1)})-\overline{g}^{(n-1)}(\widetilde{\mathbf{X}}_{p}^{(n-1)})\bigr|\right$.
$+\sum_{p=1}^{P}$ $\bigl|\overline{g}^{(n-1)}(\widetilde{\mathbf{X}}_{p}^{(n-1)})-\overline{g}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})\bigr|+\left.\sum_{p=1}^{P}\bigl|\overline{g}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})-\hat{g}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})\bigr|\right).$
For the estimate of the first part on the right-hand side of (59), note that $\widetilde{\mathbf{X}}_{p}^{(n-1)}$ are i.i.d. and follow the same distribution as $\mathbf{X}^{(n-1)}$. Its expectation is given by
(60) $\quad\mathbb{E}\bigl|g^{(n-1)}(\widehat{\mathbf{X}}^{(n-1)})-\overline{g}^{(n-1)}(\widehat{\mathbf{X}}^{(n-1)})\bigr|$
$=\frac{\chi\tau}{M_{0}}\int_{\Omega}\left|\nabla c^{(n-1)}(\mathbf{x})-\nabla\hat{c}^{(n-1)}(\mathbf{x})\right|\rho^{(n-1)}\mathrm{d}\mathbf{x}$
$\leq\frac{\chi\tau}{M_{0}}\left(\int_{\Omega}\left|\nabla c^{(n-1)}(\mathbf{x})-\nabla\hat{c}^{(n-1)}(\mathbf{x})\right|^{2}\mathrm{d}\mathbf{x}\right)^{\frac{1}{2}}\left(\int_{\Omega}\left(\rho^{(n-1)}\right)^{2}\mathrm{d}\mathbf{x}\right)^{\frac{1}{2}}$
$\leq\frac{\chi\tau}{M_{0}}\left(\int_{\Omega}\left|\nabla c^{(n-1)}(\mathbf{x})-\nabla\hat{c}^{(n-1)}(\mathbf{x})\right|^{2}\mathrm{d}\mathbf{x}\right)^{\frac{1}{2}}N_{2}^{\frac{1}{2}},$
and its variance has an upper bound
(61) $\mathrm{Var}\bigl|g^{(n-1)}(\widehat{\mathbf{X}}^{(n-1)})$ $-\overline{g}^{(n-1)}(\widehat{\mathbf{X}}^{(n-1)})\bigr|\leq\mathbb{E}\bigl|g^{(n-1)}(\widehat{\mathbf{X}}^{(n-1)})-\overline{g}^{(n-1)}(\widehat{\mathbf{X}}^{(n-1)})\bigr|^{2}$
$=\frac{\chi^{2}\tau^{2}}{M_{0}}\int_{\Omega}\left|\nabla c^{(n-1)}(\mathbf{x})-\nabla\hat{c}^{(n-1)}(\mathbf{x})\right|^{2}\rho^{(n-1)}\mathrm{d}\mathbf{x}$
$\leq\frac{\chi^{2}\tau^{2}}{M_{0}}\int_{\Omega}4M_{1}^{2}\rho^{(n-1)}\mathrm{d}\mathbf{x}=\chi^{2}\tau^{2}\cdot 4M_{1}^{2}.$
We have the following Chernoff bound
(62) $\quad\frac{1}{P}\sum_{p=1}^{P}\bigl|g^{(n-1)}(\widetilde{\mathbf{X}}_{p}^{(n-1)})-\overline{g}^{(n-1)}(\widetilde{\mathbf{X}}_{p}^{(n-1)})\bigr|$
$\leq\mathbb{E}\bigl|g^{(n-1)}(\mathbf{X}^{(n-1)})-\overline{g}^{(n-1)}(\mathbf{X}^{(n-1)})\bigr|+\frac{S_{1}}{P^{\frac{1}{2}}}\mathrm{Var}\bigl|g^{(n-1)}(\mathbf{X}^{(n-1)})-\overline{g}^{(n-1)}(\mathbf{X}^{(n-1)})\bigr|$
$\leq\frac{\chi\tau}{M_{0}}\left(\int_{\Omega}\left|\nabla c^{(n-1)}(\mathbf{x})-\nabla\hat{c}^{(n-1)}(\mathbf{x})\right|^{2}\mathrm{d}\mathbf{x}\right)^{\frac{1}{2}}N_{2}^{\frac{1}{2}}+\frac{S_{1}}{P^{\frac{1}{2}}}\chi\tau\cdot 2M_{1}$
with probability at least $1-e^{-\frac{S_{1}^{2}}{2}}$.
For the second part on the right-hand side of (59), we employ Assumption 4.5, which states that the second derivatives of $\hat{c}$ are bounded by $M_{2}$. Define the displacements as $\epsilon_{p}^{(n-1)}:=\widehat{\mathbf{X}}_{p}^{(n-1)}-\widetilde{\mathbf{X}}_{p}^{(n-1)}$. Recall that $d^{(n)}=\frac{1}{P}\sum\limits_{p=1}^{P}|\epsilon_{p}^{(n)}|$, and
(63) $\quad\frac{1}{P}\sum_{p=1}^{P}\left|\overline{g}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})-\overline{g}^{(n-1)}(\widetilde{\mathbf{X}}_{p}^{(n-1)})\right|$
$=\frac{1}{P}\sum_{p=1}^{P}\left|\epsilon_{p}^{(n-1)}\left(I_{3}+\chi\tau\nabla^{2}\hat{c}\left(\widetilde{\mathbf{X}}_{p}^{(n-1)}+\lambda_{p}\epsilon_{p}^{(n-1)}\right)\right)\right|$
$\leq\frac{1}{P}\sum_{p=1}^{P}\left|\epsilon_{p}^{(n-1)}\right|\left(1+\chi\tau M_{2}\right)=d^{(n-1)}\left(1+\chi\tau M_{2}\right).$
By Lemma 3.11, the last term on the right-hand side of (59) admits the following bound
(64) $\quad\frac{1}{P}\sum_{p=1}^{P}\bigl|\overline{g}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})-\hat{g}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})\bigr|$
$=\frac{\chi\tau}{P}\sum_{p=1}^{P}\left|\nabla\hat{c}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})-I^{*}\nabla\hat{c}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})\right|$
$\leq\frac{\chi\tau}{P}\sum_{p=1}^{P}M_{3}\left(\frac{L}{H}\right)^{2}=\chi\tau M_{3}\left(\frac{L}{H}\right)^{2}.$
Combining the three estimates via (62), (63), and (64), we obtain the target estimate in Lemma 4.8.
The discrepancy between the auxiliary and empirical gradient fields is estimated as follows.
Lemma 4.10.
The $2$-norm of the difference between the auxiliary gradient $\nabla\tilde{c}$ and the empirical gradient $\nabla\hat{c}$ over the domain $\Omega$ is bounded by
(65) $\int_{\Omega}\left|\nabla\tilde{c}^{(n-1)}(\mathbf{x})-\nabla\hat{c}^{(n-1)}(\mathbf{x})\right|^{2}\mathrm{d}\mathbf{x}\leq 2M_{0}^{2}T\left(D^{(n)}\right)^{2}+2\cdot(2\pi)^{6}M_{0}^{2}C^{2}L^{-1}H_{0}^{9}H^{-8},$
where $D^{(n)}=\max(d^{(0)},\ldots,d^{(n-1)})$ denotes the maximum historical error across iterations, $T$ is the total computational time, and the interpolation error constant is $C=\frac{513}{16}$ for the fourth-order particle-to-grid interpolation.
Proof 4.11.
We use Parseval’s theorem to establish the required bound:
(66) $\int_{\Omega}\left|\nabla\tilde{c}^{(n-1)}(\mathbf{x})-\nabla\hat{c}^{(n-1)}(\mathbf{x})\right|^{2}\mathrm{d}\mathbf{x}=L^{3}\sum_{\mathbf{q}\in U_{0}}\left|\mathbf{q}\frac{2\pi}{L}\cdot e^{(n)}_{\mathbf{q}}\right|^{2},$
where the truncated frequency index set is defined as $U_{0}:=\left\{-\frac{H_{0}}{2},\ldots,\frac{H_{0}}{2}\right\}^{3}.$
The dynamics of spectral error $e_{\mathbf{q}}^{(n)}$ defined in Definition 4.3 follow directly from the update formulas (10) and (41). For any frequency index $\mathbf{q}\in U_{0}$,
(67) $e^{(n)}_{\mathbf{q}}$ $=\frac{1}{1+\frac{\tau}{\epsilon}\left(\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}\right)}e^{(n-1)}_{\mathbf{q}}$
$\quad+\frac{1}{\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}+\frac{\epsilon}{\tau}}\biggl((\tilde{\beta}^{(n-1)}_{\mathbf{q}}-\hat{\beta}^{(n-1)}_{\mathbf{q}})+(\hat{\beta}^{(n-1)}_{\mathbf{q}}-\overline{\beta}^{(n-1)}_{\mathbf{q}})\biggr)$,
where $\tilde{\beta}_{\mathbf{q}}:=\mathcal{F}_{H}\tilde{\rho}(\mathbf{q}\frac{2\pi}{L})$, $\hat{\beta}_{\mathbf{q}}:=\mathcal{F}_{H}\hat{\rho}(\mathbf{q}\frac{2\pi}{L})$, and $\overline{\beta}_{\mathbf{q}}:=\mathcal{F}_{H}I\left(\hat{\rho}\right)(\mathbf{q}\frac{2\pi}{L})$ denote the discrete Fourier coefficients of the respective density fields.
By Lemma 3.9, we have the following bound for the first term:
(68) $|\hat{\beta}^{(n-1)}_{\mathbf{q}}-\overline{\beta}^{(n-1)}_{\mathbf{q}}|\leq CM_{0}\left|\mathbf{q}\frac{2\pi}{L}\right|^{4}\left(\frac{L}{H}\right)^{4}L^{-3}.$
Recall the displacement definition $\epsilon_{p}^{(n-1)}:=\widehat{\mathbf{X}}_{p}^{(n-1)}-\widetilde{\mathbf{X}}_{p}^{(n-1)}$. We have the following bound for the second term:
(69) $\bigl|\tilde{\beta}_{\mathbf{q}}^{(n-1)}-\hat{\beta}_{\mathbf{q}}^{(n-1)}\bigr|\leq\frac{M_{0}}{PL^{3}}\left|\sum_{p=1}^{P}e^{-i\widetilde{\mathbf{X}}_{p}^{(n-1)}\mathbf{q}\frac{2\pi}{L}}\epsilon_{p}^{(n-1)}\right|\left|\mathbf{q}\frac{2\pi}{L}\right|.$
We next define the spectral displacement error $\nu_{\mathbf{q}}^{(n)}:=\frac{1}{PL^{3}}\sum_{p=1}^{P}e^{-i\widetilde{\mathbf{X}}_{p}^{(n)}\mathbf{q}\frac{2\pi}{L}}\epsilon_{p}^{(n)}$, and the interpolation errors $\hat{e}_{\mathbf{q}}^{(n)}$ and displacement-induced errors $\check{e}_{\mathbf{q}}^{(n)}$ as
(70) $\hat{e}_{\mathbf{q}}^{(0)}=0,\hat{e}_{\mathbf{q}}^{(n)}=\frac{1}{1+\frac{\tau}{\epsilon}\left(\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}\right)}e^{(n-1)}_{\mathbf{q}}+\frac{1}{\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}+\frac{\epsilon}{\tau}}(\hat{\beta}^{(n-1)}_{\mathbf{q}}-\overline{\beta}^{(n-1)}_{\mathbf{q}})$,
$\check{e}_{\mathbf{q}}^{(0)}=0,\check{e}_{\mathbf{q}}^{(n)}=\frac{1}{1+\frac{\tau}{\epsilon}\left(\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}\right)}e^{(n-1)}_{\mathbf{q}}+\frac{1}{\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}+\frac{\epsilon}{\tau}}(\tilde{\beta}^{(n-1)}_{\mathbf{q}}-\hat{\beta}^{(n-1)}_{\mathbf{q}})$.
By (67) we have $|e_{\mathbf{q}}^{(n)}|\leq|\hat{e}_{\mathbf{q}}^{(n)}|+|\check{e}_{\mathbf{q}}^{(n)}|.$
These two error terms admit the following estimates, respectively
(71) $|\hat{e}_{\mathbf{q}}^{(n)}|\leq\frac{1}{1+\frac{\tau}{\epsilon}\left(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}\right)}|\hat{e}_{\mathbf{q}}^{(n-1)}|+\frac{M_{0}L^{-3}}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}+\frac{\epsilon}{\tau}}\cdot C\left|\mathbf{q}\frac{2\pi}{L}\right|^{4}\left(\frac{L}{H}\right)^{4},$
and
(72) $|\check{e}_{\mathbf{q}}^{(n)}|\leq\frac{1}{1+\frac{\tau}{\epsilon}\left(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}\right)}|\check{e}_{\mathbf{q}}^{(n-1)}|+\frac{M_{0}}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}+\frac{\epsilon}{\tau}}\cdot\left|\mathbf{q}\frac{2\pi}{L}\right|\left|\nu_{\mathbf{q}}^{(n-1)}\right|.$
For (71), summing the decaying geometric series yields
(73) $|\hat{e}_{\mathbf{q}}^{(n)}|$ $\leq\frac{1+\frac{\tau}{\epsilon}\left(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}\right)}{\frac{\tau}{\epsilon}\left(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}\right)}\cdot\frac{M_{0}L^{-3}}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}+\frac{\epsilon}{\tau}}\cdot C\left|\mathbf{q}\frac{2\pi}{L}\right|^{4}\left(\frac{L}{H}\right)^{4}$
$=\frac{M_{0}L^{-3}}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}}\cdot C\left|\mathbf{q}\frac{2\pi}{L}\right|^{4}\left(\frac{L}{H}\right)^{4}.$
For (72), we first define the single-step decay factor as $\kappa_{\mathbf{q}}=\frac{1}{1+\frac{\tau}{\epsilon}\left(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}\right)}$, and then derive the estimate as follows:
(74) $|\check{e}_{\mathbf{q}}^{(n)}|^{2}$ $\leq\frac{M_{0}^{2}\tau^{2}}{\epsilon^{2}}\left(\left|\nu_{\mathbf{q}}^{(n-1)}\right|+\kappa_{\mathbf{q}}\left|\nu_{\mathbf{q}}^{(n-2)}\right|+\cdots+\kappa_{\mathbf{q}}^{n-1}\left|\nu_{\mathbf{q}}^{(0)}\right|\right)^{2}$
$\leq\frac{M_{0}^{2}\tau^{2}}{\epsilon^{2}}\left(1+\kappa_{\mathbf{q}}^{2}+\cdots+\kappa_{\mathbf{q}}^{2n-2}\right)\left(\left|\nu_{\mathbf{q}}^{(n-1)}\right|^{2}+\left|\nu_{\mathbf{q}}^{(n-2)}\right|^{2}+\cdots+\left|\nu_{\mathbf{q}}^{(0)}\right|^{2}\right)$
$\leq\frac{M_{0}^{2}\tau^{2}}{\epsilon^{2}}\cdot\frac{1}{1-\kappa_{\mathbf{q}}^{2}}\left(\left|\nu_{\mathbf{q}}^{(n-1)}\right|^{2}+\left|\nu_{\mathbf{q}}^{(n-2)}\right|^{2}+\cdots+\left|\nu_{\mathbf{q}}^{(0)}\right|^{2}\right)$
$\leq\frac{M_{0}^{2}\tau}{\epsilon}\cdot\frac{1}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}}\left(\left|\nu_{\mathbf{q}}^{(n-1)}\right|^{2}+\left|\nu_{\mathbf{q}}^{(n-2)}\right|^{2}+\cdots+\left|\nu_{\mathbf{q}}^{(0)}\right|^{2}\right).$
Applying the above spectral error bounds and using Parseval’s identity (66), we translate the frequency-domain estimates back to the physical space to proceed with the proof. Specifically, we can obtain that
(75) $\sum_{\mathbf{q}\in U_{0}}\left|\mathbf{q}\frac{2\pi}{L}\cdot e_{\mathbf{q}}^{(n)}\right|^{2}\leq 2\sum_{\mathbf{q}\in U_{0}}\left|\mathbf{q}\frac{2\pi}{L}\cdot\hat{e}_{\mathbf{q}}^{(n)}\right|^{2}+2\sum_{\mathbf{q}\in U_{0}}\left|\mathbf{q}\frac{2\pi}{L}\cdot\check{e}_{\mathbf{q}}^{(n)}\right|^{2},$
(76) $\sum_{\mathbf{q}\in U_{0}}\left|\mathbf{q}\frac{2\pi}{L}\cdot\hat{e}_{\mathbf{q}}^{(n)}\right|^{2}$ $\leq M_{0}^{2}L^{-6}\sum_{\mathbf{q}\in U_{0}}\left(\frac{C\left|\mathbf{q}\frac{2\pi}{L}\right|^{5}}{\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}}\left(\frac{L}{H}\right)^{4}\right)^{2}$
$\leq M_{0}^{2}L^{-6}\sum_{\mathbf{q}\in U_{0}}C^{2}\left|\mathbf{q}\frac{2\pi}{L}\right|^{6}\left(\frac{L}{H}\right)^{8}$
$\leq H_{0}^{3}\cdot M_{0}^{2}C^{2}L^{-6}\left(H_{0}\frac{2\pi}{L}\right)^{6}\left(\frac{L}{H}\right)^{8},$
and
(77) $\sum_{\mathbf{q}\in U_{0}}\left|\mathbf{q}\frac{2\pi}{L}\cdot\check{e}_{\mathbf{q}}^{(n)}\right|^{2}$ $\leq\frac{M_{0}^{2}\tau}{\epsilon}\sum_{\mathbf{q}\in U_{0}}\frac{|\mathbf{q}\frac{2\pi}{L}|^{2}}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}}\sum_{s=0}^{n-1}\left|\nu_{\mathbf{q}}^{(s)}\right|^{2}$
$\leq\ \frac{M_{0}^{2}\tau}{\epsilon}\sum_{s=0}^{n-1}\sum_{\mathbf{q}\in U_{0}}\left|\nu_{\mathbf{q}}^{(s)}\right|^{2}.$
By Parseval’s theorem, the single-step displacement square norm $\sum_{\mathbf{q}\in U_{0}}\left|\nu_{\mathbf{q}}^{(s)}\right|^{2}=O\left(L^{-3}(d^{(s)})^{2}\right)$. Since the number of time steps satisfies $n\leq N_{T}=\frac{T}{\tau}$, the accumulated displacement square norm is $\sum_{0\leq s<n}\sum_{\mathbf{q}\in U_{0}}\left|\nu_{\mathbf{q}}^{(s)}\right|^{2}=O\left(N_{T}L^{-3}(D^{(n)})^{2}\right)$. Finally, substituting (75)-(77) into (66) yields the desired result.
Remark 4.12.
The spectral displacement error $\nu_{\mathbf{q}}^{(n)}=\frac{1}{PL^{3}}\sum_{p=1}^{P}e^{-i\widetilde{\mathbf{X}}_{p}^{(n)}\mathbf{q}\frac{2\pi}{L}}\epsilon_{p}^{(n)},$ is a tensor of size $H\times H\times H\times 3$, with the Fourier transform applied to its first three dimensions.
The discrepancy between the reference and auxiliary gradient fields is estimated as follows.
Lemma 4.13.
The $2$-norm of the difference between the reference gradient $\nabla c$ and the auxiliary gradient $\nabla\tilde{c}$ is bounded by
(78) $\mathbb{E}\left(\int_{\Omega}\left|\nabla c^{(n-1)}(\mathbf{x})-\nabla\tilde{c}^{(n-1)}(\mathbf{x})\right|^{2}\mathrm{d}\mathbf{x}\right)\leq\frac{M_{0}^{2}}{6P}L^{-1}H_{0}+(2\pi)^{-4}L^{4}N_{1,2}{H_{0}}^{-4},$
where $N_{1,2}$ is defined in Assumption 4.5, and the underlying probability space is specified in Remark 4.2.
Proof 4.14.
By Parseval’s theorem and noting that $\tilde{\alpha}^{(n)}_{\mathbf{q}}=0$ for $q\in\mathbb{Z}^{3}\setminus U_{0}$, we have
(79) $\int_{\Omega}\left|\nabla c^{(n-1)}(\mathbf{x})-\nabla\tilde{c}^{(n-1)}(\mathbf{x})\right|^{2}\mathrm{d}\mathbf{x}$
$=L^{3}\biggl(\sum_{\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}}\left|\mathbf{q}\frac{2\pi}{L}\alpha^{(n)}_{\mathbf{q}}\right|^{2}+\sum_{\mathbf{q}\in U_{0}}\left|\mathbf{q}\frac{2\pi}{L}(\alpha^{(n)}_{\mathbf{q}}-\tilde{\alpha}^{(n)}_{\mathbf{q}})\right|^{2}\biggr).$
For high frequencies $\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}$, the error satisfies $e^{(n)}_{\mathbf{q}}=\alpha^{(n)}_{\mathbf{q}}$. The update formula for $\alpha_{\mathbf{q}}$ is given by
(80) $\alpha^{(n)}_{\mathbf{q}}=\frac{1}{1+\frac{\tau}{\epsilon}(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2})}\alpha^{(n-1)}_{\mathbf{q}}+\frac{\beta^{(n-1)}_{\mathbf{q}}}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}+\frac{\epsilon}{\tau}}+O(\tau^{2}).$
Applying the inequality $(u+v)^{2}\leq(1+s)u^{2}+\left(1+\frac{1}{s}\right)v^{2}$ with $s=\frac{\tau}{\epsilon}\left(\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}\right)$, $u=\frac{\left|\alpha^{(n-1)}_{\mathbf{q}}\right|}{1+s}$, and $v=\frac{\left|\beta^{(n-1)}_{\mathbf{q}}\right|}{\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}+k^{2}+\frac{\epsilon}{\tau}}$, we obtain the following bound:
(81) $|\alpha^{(n)}_{\mathbf{q}}|^{2}\leq\frac{|\alpha^{(n-1)}_{\mathbf{q}}|^{2}}{1+\frac{\tau}{\epsilon}(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2})}+\frac{|\beta^{(n-1)}_{\mathbf{q}}|^{2}}{(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2})(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}+\frac{\epsilon}{\tau})}.$
For all $\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}$, multiplying through by $|\mathbf{q}\frac{2\pi}{L}|^{2}$ yields
(82) $\left|\mathbf{q}\frac{2\pi}{L}\alpha^{(n)}_{\mathbf{q}}\right|^{2}$ $\leq\frac{\left|\mathbf{q}\frac{2\pi}{L}\alpha^{(n-1)}_{\mathbf{q}}\right|^{2}}{1+\frac{\tau}{\epsilon}\left((\mathbf{q}\frac{2\pi}{L})^{2}+k^{2}\right)}+\frac{\left|\mathbf{q}\frac{2\pi}{L}\beta^{(n-1)}_{\mathbf{q}}\right|^{2}}{(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2})(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}+\frac{\epsilon}{\tau})}$
$\leq\frac{\left|\mathbf{q}\frac{2\pi}{L}\alpha^{(n-1)}_{\mathbf{q}}\right|^{2}}{1+\frac{\tau}{\epsilon}\left((H_{0}\frac{2\pi}{L})^{2}+k^{2}\right)}+\frac{\left|\beta^{(n-1)}_{\mathbf{q}}\right|^{2}}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}+\frac{\epsilon}{\tau}}.$
Summing over $\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}$ gives
(83) $\sum_{\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}}\left|\mathbf{q}\frac{2\pi}{L}\alpha^{(n)}_{\mathbf{q}}\right|^{2}\leq\sum_{\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}}\biggl(\frac{1}{1+\frac{\tau}{\epsilon}\left((H_{0}\frac{2\pi}{L})^{2}+k^{2}\right)}\left|\mathbf{q}\frac{2\pi}{L}\alpha^{(n-1)}_{\mathbf{q}}\right|^{2}+\frac{\left|\beta^{(n-1)}_{\mathbf{q}}\right|^{2}}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}+\frac{\epsilon}{\tau}}\biggr).$
Following the same argument as in (73), this simplifies to
(84) $\sum_{\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}}\left|\mathbf{q}\frac{2\pi}{L}\alpha^{(n)}_{\mathbf{q}}\right|^{2}$ $\leq\max_{0\leq m<n}\sum_{\mathbf{q}\in\mathbb{Z}^{3}\setminus U_{0}}\frac{\left|\beta^{(m)}_{\mathbf{q}}\right|^{2}}{(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2})}$
$\leq\max_{0\leq m<n}\left(H_{0}\frac{2\pi}{L}\right)^{-4}\sum_{\mathbf{q}\in\mathbb{Z}^{3}}\left|\mathbf{q}\frac{2\pi}{L}\beta^{(m)}_{\mathbf{q}}\right|^{2}$
$\leq\left(H_{0}\frac{2\pi}{L}\right)^{-4}L^{-3}N_{1,2}.$
For the low-frequency terms, the independence of the particles $\widetilde{\mathbf{X}}_{1}^{(n)},\ldots,\widetilde{\mathbf{X}}_{P}^{(n)}\sim\mathbf{X}^{(n)}$ implies that
(85) $\mathbb{E}\left|\beta^{(n)}_{\mathbf{q}}-\tilde{\beta}^{(n)}_{\mathbf{q}}\right|^{2}$ $=\frac{1}{P}\mathrm{Var}\!\left(M_{0}L^{-3}e^{-i\mathbf{q}\frac{2\pi}{L}\widehat{\mathbf{X}}^{(n)}}\right)$
$=\frac{M_{0}^{2}}{P}L^{-6}\left(1-\left|\beta^{(n)}_{2\mathbf{q}}\right|^{2}\right)$
$\leq\frac{M_{0}^{2}}{P}L^{-6}.$
Applying the attenuation factor $\frac{1}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}}$ and taking the expectation over the sum yields
(86) $\mathbb{E}\left(\sum_{\mathbf{q}\in U_{0}}\left|\mathbf{q}\frac{2\pi}{L}(\alpha^{(n)}_{\mathbf{q}}-\tilde{\alpha}^{(n)}_{\mathbf{q}})\right|^{2}\right)$ $\leq\sum_{\mathbf{q}\in U_{0}}\frac{|\mathbf{q}\frac{2\pi}{L}|^{2}}{(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2})^{2}}\cdot\frac{M_{0}^{2}}{P}L^{-6}$
$\leq\frac{6M_{0}^{2}H_{0}}{4\pi^{2}P}L^{-4}.$
The factor $6H_{0}$ arises from a spherical shell summation: for each $1\leq h\leq H_{0}$, there are $6h^{2}$ frequency indices $\mathbf{q}$ with $L^{1}$-norm $|\mathbf{q}|_{1}=h$. The $h^{-2}$ scaling of the spectral terms $\frac{|\mathbf{q}\frac{2\pi}{L}|^{2}}{(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2})^{2}}$ cancels this growth, leaving a linear accumulation in $H_{0}$.
Combining the high-frequency error bound (84) and the low-frequency variance estimate (86) via the Parseval decomposition (79), we establish the desired final error bound.
The temporal error of the Euler-Maruyama semi-discretization over the time interval is stated as follows (see [milstein2004stochastic, p. 15-16]).
Lemma 4.15.
Let $\{\mathbf{X}_{p}\}_{1\leq p\leq P}$ be $P$ independent realizations of the SDE (11), and let the Brownian increments $\mathbf{W}_{p}$ satisfy
(87) $\mathbf{W}_{p}(n\tau)-\mathbf{W}_{p}((n-1)\tau)=\mathbf{W}_{p}^{(n)}.$
The Euler-Maruyama semi-discretization exhibits first-order strong convergence for additive noise, i.e., when the diffusion term is independent of the solution process. More precisely, for $t=n\tau\leq T$ with integer $n\geq 1$,
(88) $\bigl(\mathbb{E}\bigl[|\mathbf{X}_{p}(t)-\widetilde{\mathbf{X}}_{p}^{(n)}|^{2}\bigr]\bigr)^{\frac{1}{2}}\leq C_{0}(t+\sqrt{t})\tau,$
where $C_{0}=C_{0}(J_{1},M_{1},M_{2},M_{3})$ is a positive constant independent of the time step $\tau$ and $t$.
The main convergence result is presented as follows.
Theorem 4.16.
For the SIPF-PIC method with fourth-order particle-to-grid and second-order grid-to-particle interpolation, the mean absolute error at the final time T, defined as $\hat{d}^{(n)}:=\frac{1}{P}\sum_{1\leq p\leq P}|\mathbf{X}_{p}(n\tau)-\widehat{\mathbf{X}}_{p}^{(n)}|$, satisfies the following growth bound:
[…]
$+\left.\left.C_{3}L^{2}H_{0}^{-2}+M_{3}L^{2}H^{-2}+2S_{1}M_{1}P^{-\frac{1}{2}}\right)\right),$
with probability at least $1-N_{T}e^{-\frac{S_{1}^{2}}{2}}$. Here, $\chi$ denotes the chemotactic sensitivity in the KS system (1), and $K$, $C_{0}$, $C_{1}$, $C_{2}$, $C_{3}$, and $S_{1}$ are positive constants derived in the proof, independent of the domain size $L$, grid resolution $H$, frequency cutoff threshold $H_{0}$, number of particles $P$, and time step $\tau$.
Proof 4.17.
First, we employ the discrete Gronwall inequality (Lemma 4.6) to derive an estimate for $d^{(n)}$. From Lemma 4.8, we have
(90) $F^{(n-1)}\leq M_{3}\left(\frac{L}{H}\right)^{2}+2P^{-\frac{1}{2}}S_{1}M_{1}+\frac{N_{2}^{\frac{1}{2}}}{M_{0}}\left(\int_{\Omega}\left|\nabla c^{(n-1)}(\mathbf{x})-\nabla\hat{c}^{(n-1)}(\mathbf{x})\right|^{2}\mathrm{d}\mathbf{x}\right)^{\frac{1}{2}}.$
We decompose the gradient error term using the auxiliary-empirical and reference-auxiliary bounds from Lemmas 4.10 and 4.13, respectively. The gradient discrepancy is first split via the inequality as follows:
(91) $\quad\int_{\Omega}\left|\nabla c^{(n-1)}(\mathbf{x})-\nabla\hat{c}^{(n-1)}(\mathbf{x})\right|^{2}\mathrm{d}\mathbf{x}$
$\leq 2\int_{\Omega}\left(\left|\nabla\tilde{c}^{(n-1)}(\mathbf{x})-\nabla\hat{c}^{(n-1)}(\mathbf{x})\right|^{2}+\left|\nabla c^{(n-1)}(\mathbf{x})-\nabla\tilde{c}^{(n-1)}(\mathbf{x})\right|^{2}\right)\mathrm{d}\mathbf{x},$
Substituting the reference-auxiliary expectation bound (78) and the auxiliary-empirical bound (65) into the Gronwall framework yields the composite error growth constant $K=\sqrt{2TN_{2}}$ and residual term
(92) $G^{(n)}\leq C_{1}L^{-\frac{1}{2}}H^{-4}H_{0}^{\frac{9}{2}}+C_{2}L^{-\frac{1}{2}}P^{-\frac{1}{2}}H_{0}^{\frac{1}{2}}+C_{3}L^{2}H_{0}^{-2}+M_{3}L^{2}H^{-2}+2S_{1}M_{1}P^{-\frac{1}{2}},$
where the constants $C_{1}=513\pi^{3}\sqrt{N_{2}}$, $C_{2}=\sqrt{N_{2}}$, and $C_{3}=(2\pi)^{-2}\frac{\sqrt{N_{2}N_{1,2}}}{M_{0}}$ are defined in terms of the problem-dependent constants. The final error bound for $d^{\left(N_{T}\right)}$ follows from the estimate $(1+(K+\chi M_{2})\tau)^{N_{T}}\leq e^{T(K+\chi M_{2})}$ and the temporal accumulation of the residual terms, leading to the estimate as follows:
[…]
$+\left.\left.C_{3}L^{2}H_{0}^{-2}+M_{3}L^{2}H^{-2}+2S_{1}M_{1}P^{-\frac{1}{2}}\right)\right).$
Finally, from the estimate
(94) $\hat{d}^{(N_{T})}\leq d^{(N_{T})}+\frac{1}{P}\sum_{p=1}^{P}\bigl|\mathbf{X}_{p}(T)-\widetilde{\mathbf{X}}_{p}^{(N_{T})}\bigr|\leq d^{(N_{T})}+\Biggl(\frac{1}{P}\sum_{p=1}^{P}\bigl|\mathbf{X}_{p}(T)-\widetilde{\mathbf{X}}_{p}^{(N_{T})}\bigr|^{2}\Biggr)^{\!\frac{1}{2}},$
which follows from the triangle inequality and the Cauchy-Schwarz inequality, combining with the temporal error bound from Lemma 4.15, we obtain the claimed convergence bound in (89).
Corollary 4.18.
By optimizing the frequency cutoff threshold $H_{0}=H^{\frac{8}{13}}$, the error bound simplifies to
[…]
$+\left.\left.C_{2}L^{-\frac{1}{2}}P^{-\frac{1}{2}}H^{\frac{4}{13}}+M_{3}L^{2}H^{-2}+2S_{1}M_{1}P^{-\frac{1}{2}}\right)\right)$
with probability at least $1-N_{T}e^{-\frac{S_{1}^{2}}{2}}$.
4.2 Relaxation of the smoothness assumption
In this section, we establish several conclusions showing that under specific parameter combinations, the smoothness of $\hat{c}$ in Assumption 4.5 can be naturally guaranteed by the algorithm itself.
Lemma 4.19.
Assume the particle number satisfies $P\geq H^{\frac{40}{13}}$ and the frequency cutoff threshold is set to $H_{0}=H^{\frac{8}{13}}$. Then, the condition (50) in Assumption 4.5 can be relaxed to:
(96) $\|D_{\mathbf{k}}c\|_{\infty}$ $\leq M_{\eta},$
$\|D_{\mathbf{k}}\hat{c}\|_{\infty}$ $\leq Q_{\eta}\|D_{\mathbf{k}}\hat{c}\|_{2}+R_{\eta},$
for any multi-index $\mathbf{k}$ with $|\mathbf{k}|=\eta$, where $\eta=1,2,3$, and for some constants $M_{\eta},Q_{\eta},R_{\eta}$ that are independent of $H$, $H_{0}$, and $P$.
Proof 4.20.
We show that $\|D_{\mathbf{k}}\hat{c}\|_{2}$ is controlled by $\|D_{\mathbf{k}}c\|_{2}$. For $\eta=1,2,3$, by Parseval’s theorem,
(97) $\sum_{|\mathbf{k}|=\eta}\int_{\Omega}|D_{\mathbf{k}}\hat{c}^{(n)}(\mathbf{x})|^{2}\,\mathrm{d}\mathbf{x}$ $=L^{3}\sum_{|\mathbf{k}|=\eta}\sum_{\mathbf{q}\in U}\left(\mathbf{q}\frac{2\pi}{L}\right)^{2\mathbf{k}}\left|\hat{\alpha}^{(n)}_{\mathbf{q}}\right|^{2}$
$\leq 2L^{3}\sum_{|\mathbf{k}|=\eta}\sum_{\mathbf{q}\in U}\left(\mathbf{q}\frac{2\pi}{L}\right)^{2\mathbf{k}}\left(\left|\alpha^{(n)}_{\mathbf{q}}\right|^{2}+\left|\hat{\alpha}^{(n)}_{\mathbf{q}}-\alpha^{(n)}_{\mathbf{q}}\right|^{2}\right)$
[…]
By Lemmas 4.10 and 4.13, we can obtain the estimate
(98) $\quad L^{3}\left(H_{0}\frac{2\pi}{L}\right)^{2\eta-2}\sum_{\mathbf{q}\in U}\left|\mathbf{q}\frac{2\pi}{L}\right|^{2}\left|\hat{\alpha}^{(n)}_{\mathbf{q}}-\alpha^{(n)}_{\mathbf{q}}\right|^{2}$
$\leq\left(\frac{2\pi}{L}\right)^{2\eta-2}\left(C_{4}H_{0}^{2\eta+7}H^{-8}+C_{5}H_{0}^{2\eta-1}P^{-1}+C_{6}H_{0}^{2\eta-6}+C_{7}(D^{(n)})^{2}H_{0}^{2\eta-2}\right),$
where the constants $C_{4}=M_{0}^{2}N_{2}^{-1}C_{1}^{2}L^{-1}$, $C_{5}=M_{0}^{2}N_{2}^{-1}C_{2}^{2}L^{-1}$, $C_{6}=L^{4}(2\pi)^{-4}N_{1,2}$, and $C_{7}=2M_{0}^{2}T$ are independent of $H$, $H_{0}$, $P$, and $M_{\eta}$.
For the parameters $P\geq H^{\frac{40}{13}}$ and $H_{0}=H^{\frac{8}{13}}$, we have $\|D_{\mathbf{k}}\hat{c}^{(n)}\|_{2}\leq 2\|D_{\mathbf{k}}c^{(n)}\|_{2}+B_{\eta}$ for $D^{(n)}\leq H^{-\frac{16}{13}}$, for some $B_{\eta}$ independent of $H$, $H_{0}$, $P$, and $M_{\eta}$ with $\eta=1,2,3$.
[…]
Corollary 4.21.
Assume the particle number satisfies $P\geq H^{\frac{64}{13}}$ and the frequency cutoff threshold is set to $H_{0}=H^{\frac{8}{13}}$. For a timestep converging to zero more slowly, specifically, $H_{\tau}:=\tau^{-\frac{1-\sigma}{2}}<H_{0}$ for some $\sigma>0$, the condition (50) in Assumption 4.5 can be relaxed to
(99) $\|D_{\mathbf{k}}c\|_{\infty}$ $\leq M_{\eta}$
for any multi-index $\mathbf{k}$ with $|\mathbf{k}|=\eta$, where $\eta=1,2,3$.
Proof 4.22.
We verify that the conditions of Lemma 4.19 are satisfied. Let $\hat{c}^{(n)}_{\text{low}}=\sum_{\|\mathbf{q}\|\leq H_{\tau}}\hat{\alpha}_{\mathbf{q}}^{(n)}e^{-i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}$ and $\hat{c}^{(n)}_{\text{high}}=\sum_{H_{\tau}<\|\mathbf{q}\|\leq H_{0}}\hat{\alpha}_{\mathbf{q}}^{(n)}e^{-i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}$. Then, $\|D_{\mathbf{k}}\hat{c}^{(n)}\|_{\infty}\leq\\ \|D_{\mathbf{k}}\hat{c}^{(n)}_{\text{low}}\|_{\infty}+\|D_{\mathbf{k}}\hat{c}^{(n)}_{\text{high}}\|_{\infty}$. For the low-frequency part, by the Parseval Theorem,
(100) $\|D_{\mathbf{k}}\hat{c}^{(n)}_{\text{low}}\|_{\infty}$ $\leq\sum_{\|\mathbf{q}\|\leq H_{\tau}}\left|\left(\mathbf{q}\frac{2\pi}{L}\right)^{\mathbf{k}}\right||\hat{\alpha}_{\mathbf{q}}^{(n)}|$
$\leq H_{\tau}^{\frac{3}{2}}\left(\sum_{\|\mathbf{q}\|\leq H_{\tau}}\left(\mathbf{q}\frac{2\pi}{L}\right)^{2\mathbf{k}}\left|\hat{\alpha}_{\mathbf{q}}^{(n)}\right|^{2}\right)^{\frac{1}{2}}\leq H_{\tau}^{\frac{3}{2}}L^{-\frac{3}{2}}\|D_{\mathbf{k}}\hat{c}^{(n)}\|_{2}.$
For the high-frequency part, we verify that $\left|\hat{\alpha}_{\mathbf{q}}^{(n)}\right|$ (for $H_{\tau}<\|\mathbf{q}\|\leq H_{0}$) is bounded by the diffusion term $\mathbf{W}_{p}^{(n)}\sim\mathcal{N}(0,2\mu\tau I_{3})$ at each step. Recall that
(101) $\hat{\beta}_{\mathbf{q}}^{(n)}=\frac{M_{0}}{P}L^{-3}\sum_{p=1}^{P}e^{-i\widehat{\mathbf{X}}_{p}^{(n)}\cdot\frac{2\pi}{L}\mathbf{q}}=\frac{M_{0}}{P}L^{-3}\sum_{p=1}^{P}e^{-i(\mathbf{Z}_{p}^{(n)}+\mathbf{W}_{p}^{(n)})\cdot\frac{2\pi}{L}\mathbf{q}}.$
Let the summation term for the $p$-th particle be $\hat{\beta}_{\mathbf{q},p}^{(n)}=e^{-i\mathbf{X}_{p}^{(n)}\cdot\frac{2\pi}{L}\mathbf{q}}$. Then, we can obtain that $\left|\mathbb{E}(\hat{\beta}_{\mathbf{q},p}^{(n)})\right|=e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}$, and $\text{Var}(\hat{\beta}_{\mathbf{q},p}^{(n)})=1-e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}\leq 1$. Since $\hat{\beta}_{\mathbf{q},p}^{(n)}$ are independent and $\mathbb{E}(\left|\hat{\beta}_{\mathbf{q},p}^{(n)}\right|^{\xi})=1$ for all $\xi>0$, by Lindeberg’s condition, we have $\left|\hat{\beta}_{\mathbf{q}}^{(n)}\right|\leq\frac{M_{0}}{L^{3}}(e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}+\frac{S_{1}}{\sqrt{P}})$ with probability $1-e^{-\frac{S_{1}^{2}}{2}}$.
Assume that this bound holds for all $q\in U_{0}\setminus U_{\tau}$ and for $1\leq n\leq N_{T}$, with probability at least $1-H_{0}^{3}N_{T}e^{-\frac{S_{1}^{2}}{2}}$. Similar to (80), the update rule for $\hat{\alpha}_{\mathbf{q}}$ satisfies
(102) $\hat{\alpha}^{(n)}_{\mathbf{q}}=\frac{1}{1+\frac{\tau}{\epsilon}(|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2})}\hat{\alpha}^{(n-1)}_{\mathbf{q}}+\frac{\hat{\beta}^{(n-1)}_{\mathbf{q}}+b^{(n-1)}_{\mathbf{q}}}{|\mathbf{q}\frac{2\pi}{L}|^{2}+k^{2}+\frac{\epsilon}{\tau}},$
where $b^{(n-1)}_{\mathbf{q}}\leq\frac{M_{0}}{L^{3}}CH^{-4}$ is the interpolation error. Similar to (73), this accumulates to
(103) $\left|\left(\mathbf{q}\frac{2\pi}{L}\right)^{\mathbf{k}}\right||\hat{\alpha}_{\mathbf{q}}^{(n)}|\leq\left|\mathbf{q}\frac{2\pi}{L}\right|^{\eta-2}\left(\frac{M_{0}}{L^{3}}\left(e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}+\frac{S_{1}}{\sqrt{P}}+CH^{-4}\right)\right).$
Summing over all coefficients, we obtain
(104) $\|D_{\mathbf{k}}\hat{c}^{(n)}_{\text{high}}\|_{\infty}$ $\leq\sum_{H_{\tau}<\|\mathbf{q}\|\leq H_{0}}\left|\left(\mathbf{q}\frac{2\pi}{L}\right)^{\mathbf{k}}\right||\hat{\alpha}_{\mathbf{q}}^{(n)}|$
$\leq\sum_{H_{\tau}<\|\mathbf{q}\|\leq H_{0}}\left|\mathbf{q}\frac{2\pi}{L}\right|^{\eta-2}\frac{M_{0}}{L^{3}}\left(e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}+\frac{S_{1}}{\sqrt{P}}+CH^{-4}\right).$
Since $|\mathbf{q}|^{2}\tau>|\mathbf{q}|^{\frac{2\sigma}{1-\sigma}}$ for $|\mathbf{q}|>H_{\tau}$, the first term $\sum_{\|\mathbf{q}\|>H_{\tau}}\left|\mathbf{q}\frac{2\pi}{L}\right|^{\eta-2}\frac{M_{0}}{L^{3}}e^{-\mu\tau\left(\frac{2\pi}{L}\mathbf{q}\right)^{2}}\\ \leq\frac{M_{0}}{L^{3}}C_{\sigma}$ for some constant $C_{\sigma}>0$, as it tends to zero as $H_{\tau}\to\infty$. Thus,
(105) $\|D_{\mathbf{k}}\hat{c}^{(n)}_{\text{high}}\|_{\infty}$ $\leq\frac{M_{0}}{L^{3}}C_{\sigma}+(H_{0}^{3}-H_{\tau}^{3})H_{0}\frac{2\pi M_{0}}{L^{4}}\left(\frac{S_{1}}{\sqrt{P}}+CH^{-4}\right)$
$\leq\frac{M_{0}}{L^{3}}C_{\sigma}+H_{0}^{4}\frac{2\pi M_{0}}{L^{4}}\left(\frac{S_{1}}{\sqrt{P}}+CH^{-4}\right).$
From (100) and (105), under the conditions $H_{0}=H^{\frac{8}{13}}$ and $P\geq H_{0}^{8}$, we conclude the proof.
Remark 4.23.
The presented error analysis critically relies on the non-blow-up assumptions embedded in the framework. In scenarios where solutions develop singularities or lose regularity (e.g., near a delta-type blow-up), the square-integrability condition and the smoothness requirement in Assumption 4.5 no longer hold.
5 Numerical experiments
This section presents numerical results to validate the convergence analysis and demonstrate the performance of the SIPF-PIC method. We first establish convergence properties through systematic analysis of particle-grid resolution dependencies, followed by demonstrations of the framework’s capacity to resolve finite-time blow-up dynamics.
The section is organized as follows. Subsection 5.1 quantifies algorithmic convergence through error analysis with respect to the particle number $P$ and grid resolution $H$ per dimension, and subsequently compares computational performance using identical Keller-Segel parameters. Building on these foundations, Subsection 5.2 verifies the theoretical threshold $M_{\text{critical}}=8\pi$ in 2D elliptic systems and identifies an empirical threshold in 3D parabolic systems. We then explore blow-up dynamics: Subsection 5.3 reveals three mass-dependent phases in tetrahedral symmetric systems, including subcritical dispersion, intermediate unified blowup, and supercritical multistage singularity formation. Finally, Subsection 5.4 demonstrates radial concentration into singular ring structures that exhibit post-collapse instabilities.
5.1 Convergence of the SIPF-PIC method with respect to discretization
We first validate the convergence of the algorithm with respect to the particle number $P$ and grid resolution $H$ per dimension. Consider a spherically symmetric initial configuration where $\rho_{0}(\mathbf{x})$ is uniformly distributed within a unit ball centered at the origin:
(106) $\rho_{0}(\mathbf{x})=\begin{cases}\frac{60}{\pi}&|\mathbf{x}|\leq 1,\\ 0&\text{otherwise},\end{cases}$
with total mass $M_{0}=80$. The model parameters in (1) are set as $\mu=\chi=1$, $\epsilon=10^{-4}$, $k=10^{-1}$, with final time $T=2\times 10^{-2}$, timestep $\tau=10^{-5}$, and total time steps $N_{T}=2000$.
By exploiting radial symmetry, we can reduce the 3D system to a 1D radial coordinate $r$ through the following transformation
(107) $x_{1}=r\cos\theta,\quad x_{2}=r\sin\theta\cos\phi,\quad x_{3}=r\sin\theta\sin\phi.$
This yields the simplified system:
(108) $\begin{cases}\displaystyle\rho_{t}=\frac{2\mu}{r}\rho_{r}+\mu\rho_{rr}-\frac{2\chi}{r}\rho c_{r}-\chi\rho_{r}c_{r}-\chi\rho c_{rr},\\ \displaystyle\epsilon c_{t}=\frac{2}{r}c_{r}+c_{rr}-k^{2}c+\rho,\end{cases}$
together with boundary conditions $\rho_{r}|_{r=0}=c_{r}|_{r=0}=0$.
While the problem defined in (106) is posed in $\mathbb{R}^{3}$, the particle dynamics remain localized within $|\mathbf{x}|<2$ during the entire simulation. The periodic setting in $\Omega=[-10,10]^{3}$ (i.e., $L=20$) thus suffices to approximate the unbounded problem, as there is almost no density distribution outside of the domain $\Omega$.
A reference solution can then be computed using an implicit finite difference method to solve (108) on a high-resolution 1D grid with $N=5\times 10^{5}$ points, utilizing the radial symmetry.
Convergence studies for the SIPF-PIC method are conducted by varying $P$ and $H$ while comparing the numerical results with this reference solution. We use the cumulative distribution function
(109) $m(r,t;\rho)=\frac{1}{M_{0}}\int_{|\mathbf{x}|\leq r}\rho(\mathbf{x},t)\,\mathrm{d}\mathbf{x}$
at the final time $T=2\times 10^{-2}$ as the comparative criterion. Since $m(0,T;\rho)=0$ and $m(\infty,T;\rho)=1$ for every mass-preserving density function $\rho$, we compute the $W^{1}$ distance between the results obtained by the SIPF-PIC method and the reference solution as the error metric. Specifically, for the density $\hat{\rho}$ defined by (6), we have $m(r,T,\hat{\rho})=\frac{1}{P}\sum_{p=1}^{P}\mathbf{1}_{\{|X_{p}^{(N_{T})}|\leq r\}},$ and for a radially symmetric density $\rho(\mathbf{x},t)=\psi(|\mathbf{x}|,t)$, we have $m(r,T,\rho)=\frac{4\pi}{M_{0}}\int_{0}^{r}\psi(s,t)s^{2}\,ds.$ To compute the Wasserstein distance, we obtain the quantiles $q_{\frac{j}{n}}$ and $\hat{q}_{\frac{j}{n}}$ for $n=10^{4}$ and $j=0,1,\ldots,n-1$, restricted to the interval $[0,R]$, where $R=50$ is a sufficiently large radius. Specifically, we have $m(q^{*}_{\frac{j}{n}},T,\rho)=\frac{j}{n}$ and $m(\hat{q}^{*}_{\frac{j}{n}},T,\hat{\rho})=\frac{j}{n}$ with $q_{\frac{j}{n}}=\min(q^{*}_{\frac{j}{n}},R)$ and $\hat{q}_{\frac{j}{n}}=\min(\hat{q}^{*}_{\frac{j}{n}},R)$.
Finally, we define the Wasserstein $W^{1}$ error as
(110) $\varepsilon=\frac{1}{n}\sum_{j=0}^{n-1}|q_{\frac{j}{n}}-\hat{q}_{\frac{j}{n}}|.$
We test all combinations of $P\in\{2^{11},\ldots,2^{20}\}$ and $H\in\{8,16,32,64,128,256\}$. In these simulations, we set $H_{0}=H$, meaning that no frequency cutoff is applied.
Figure 3 demonstrates the convergence behavior of the SIPF-PIC method using fourth-order particle-to-grid and second-order grid-to-particle interpolations, quantifying convergence rates with respect to the individual parameters. Two complementary studies isolate the parametric dependencies: fixing the grid resolution at $H_{\max}$ yields the particle number convergence rate via the slope of $\log_{2}\mathcal{E}$ versus $\log_{2}P$, with a measured value of 0.4207-0.4207, which closely matches the theoretical order of $1/2$ derived in Subsection 4; fixing $P_{\max}$ determines the spatial convergence order through the slope of $\log_{2}\mathcal{E}$ versus $\log_{2}H$, with a measured value of 1.4975-1.4975, which closely matches the theoretical order of $16/13$.
(a) Error convergence against particle number ($P$) with fixed grid resolution.
(b) Error convergence against grid resolution ($H$) with fixed particle number.
Figure 3: Convergence characteristics under the dual parameters of particle number and grid resolution.
In addition to the convergence results presented in Figure 3, we investigate a performance comparison between the original SIPF method and our new SIPF-PIC method (using second- or fourth-order interpolation) through timing and error measurements for the KS system, using identical parameter sets and initial conditions to those in Section 5.1. The test cases employ grid resolutions $H\in\{8,24,64,256\}$ and particle numbers $P\in\{2^{11},2^{14},2^{17},2^{20}\}$, with results tabulated in Table 1.
Table 1: Time-error profiles of SIPF and SIPF-PIC with different interpolation orders.
$H$ $P$ SIPF SIPF-PIC (2/2) SIPF-PIC (4/4)
Time (s) Error Time (s) Error Time (s) Error
8 $2^{11}$ 74.4 $7.8\times 10^{-2}$ 2.2 $1.0\times 10^{-1}$ 5.3 $9.9\times 10^{-2}$
8 $2^{14}$ 2617 $6.2\times 10^{-2}$ 5.4 $8.8\times 10^{-2}$ 19.6 $8.5\times 10^{-2}$
24 $2^{14}$ 10297 $1.5\times 10^{-2}$ 7.8 $4.6\times 10^{-2}$ 21.6 $2.6\times 10^{-2}$
64 $2^{17}$ 52.4 $9.8\times 10^{-3}$ 112 $2.0\times 10^{-3}$
256 $2^{20}$ 1756 $5.2\times 10^{-4}$ 2238 $3.6\times 10^{-4}$
The complexity reduction from $O(PH^{3})$ to $O(P+H^{3}\log H)$ through the SIPF-PIC method’s localized particle-grid interpolations and spectral field solutions fundamentally decouples particle-field interactions, achieving accelerated parameter exploration while preserving numerical accuracy and avoiding the reintroduction of $PH^{3}$ bottlenecks. This complexity separation permits high-resolution simulations at extreme resolutions (e.g., $H\geq 256$ and $P\geq 2^{20}$) with controlled error growth, overcoming prior SIPF limitations via adaptive accuracy-efficiency tradeoffs.
We also demonstrate that the algorithm is consistent with respect to $H$. When the number of particles $P$ is fixed and the grid resolution $H$ is increased, the number of grid points $H^{3}$ may become significantly larger than $P$. In this scenario, according to Theorem 4.16, our algorithm still yields convergent results. We fix $P=10^{4}$ (consistent with the original SIPF method) and set the truncation frequency $H_{0}=16$, while increasing $H$ from $16$ to $256$. The error of each configuration compared to the reference solution is shown in Fig. 4. As $H$ increases, the error converges to a constant value determined by $P$ and $H_{0}$.
Figure 4: Convergence with increasing grid resolution $H$ under a fixed particle number ($P=10^{4}$) and truncation frequency $H_{0}=16$.
5.2 Critical mass in 2D and 3D systems
Recall the 2D parabolic-elliptic KS system with parameters $(k,\epsilon,\mu,\chi)=(0,0,1,1)$, where the critical mass threshold $M_{\text{crit}}=8\pi$ governs finite-time blow-up (see Section 2.1). Numerical simulations initialize with uniform distributions on the unit disk $\rho_{0}(\mathbf{x})=\frac{M_{0}}{\pi}\mathbf{1}_{|\mathbf{x}|\leq 1}$, for masses $M_{0}\in\{22.0,23.0,\ldots,27.0\}$, discretized in the spatial domain $\Omega=[-4,4]^{2}$ with grid resolution $H=2048$ on each axis and $P=2^{20}$ particles, utilizing fourth-order particle-to-grid and second-order grid-to-particle interpolations, coupled with a time step $\tau=5\times 10^{-4}$ up to final time $T=2.0$, totaling $N_{T}=4000$ steps.
Figure 5 demonstrates the particle distributions at time $T=2.0$ for different initial masses $M_{0}=22.0,23.0,\ldots,27.0$, exhibiting quantitative agreement with theoretical predictions: subcritical masses ($M_{0}<8\pi$) disperse radially, near-critical masses ($M_{0}\approx 8\pi$) maintain quasi-stationary profiles, and supercritical masses ($M_{0}>8\pi$) undergo concentration toward the origin.
Figure 5: Two-dimensional particle distributions at $T=2.0$ for different initial masses. (a)-(d): Subcritical to near-critical regimes ($M_{0}=22.0$-$25.0$); (e)-(f): Supercritical regimes ($M_{0}=26.0$-$27.0$).
To assess our numerical accuracy, we consider a series of near-critical initial masses $M_{0}=24.6,24.8,\ldots,25.6$, which are close to the theoretical critical threshold $8\pi\approx 25.1$. For an unbounded domain, the exact second moment at final time $T=2$ is given by
(111) $m_{2}(2)=\int_{\mathbb{R}^{2}}|\mathbf{x}|^{2}\rho(\mathbf{x},2)\,\mathrm{d}\mathbf{x}=\frac{17}{2}-\frac{M_{0}}{\pi}$
according to equation (3).
To approximate the unbounded domain, we simulate the system on periodic square domains of increasing sizes $L=10,20,40$, while keeping a fixed spatial mesh size $\Delta x=5/512$. The numerical second moment is computed from the particles as
(112) $\widehat{m}_{2}(2)=\frac{1}{P}\sum_{j=1}^{P}|\widehat{\mathbf{X}}_{j}(2)|^{2},$
where $P=2^{20}$ particles are used.
Figure 6: Convergence of the numerical second moment to the theoretical value. Each polyline corresponds to a different period $L$, while the dashed line represents the theoretical value.
As shown in Figure 6, the numerical second moment converges to the theoretical prediction as $L$ grows, confirming that periodic boundary effects diminish and the simulation approaches the unbounded-domain result. This agreement validates the accuracy of our numerical scheme, even near the critical mass where the dynamics are sensitive.
For the 3D system (1) with parameters $(k,\epsilon,\mu,\chi)=(10^{-1},10^{-4},1,1)$, no explicit critical mass threshold is theoretically established. Numerical experiments are conducted with initial masses $M_{0}\in\{45,46,\ldots,50\}$ uniformly distributed within the unit ball $\rho_{0}(\mathbf{x})=\frac{3M_{0}}{4\pi}\mathbf{1}_{\|\mathbf{x}\|\leq 1}$. The spatial domain $\Omega=[-4,4]^{3}$ is discretized using $H=256$ grid points per axis and $P=2^{20}$ particles.
The computational framework combines fourth-order particle-to-grid and second-order grid-to-particle interpolations, with time integration performed at a fixed time step $\tau=5\times 10^{-4}$ up to final time $T=1.0$, totaling $N_{T}=2000$ steps. The particle distributions at the final time corresponding to different initial masses are shown in Figure 7.
Figure 7 indicates a critical mass threshold between $M_{0}=48.0$ and $M_{0}=49.0$, consistent with the results of finite-difference solutions to the spherically symmetric 1D problem.
Figure 7: Three-dimensional particle distributions at $T=1.0$ for different initial masses. (a)-(d): Subcritical to near-critical regimes ($M_{0}=45.0$-$48.0$); (e)-(f): Supercritical regimes ($M_{0}=49.0$-$50.0$).
The original SIPF method [SIPF1] demonstrates that while blow-up occurs, the computed $\|\hat{c}\|_{\infty}$ grows with the truncation frequency $H_{0}$. Figure 8 shows this indicator for 2D and 3D cases, confirming critical masses consistent with those derived from particle distributions in Figures 5 and 7.
Figure 8: $\frac{\|\hat{c}\|_{\infty,H_{0}=32}}{\|\hat{c}\|_{\infty,H_{0}=8}}$ vs. computational time $T$.
5.3 Blowup dynamics in tetrahedral configuration
We investigate singularity formation using a non-radially symmetric initial configuration: four uniform density balls centered at the vertices of a regular tetrahedron (Fig. 9). Each ball has mass $\frac{M_{0}}{4}$, radius $\frac{1}{2}$ and centers at:
$\mathbf{x}_{1}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix},\;\mathbf{x}_{2}=\begin{pmatrix}-\frac{1}{2}\\ \frac{\sqrt{3}}{2}\\ 0\end{pmatrix},\;\mathbf{x}_{3}=\begin{pmatrix}-\frac{1}{2}\\ -\frac{\sqrt{3}}{2}\\ 0\end{pmatrix},\;\mathbf{x}_{4}=\begin{pmatrix}0\\ 0\\ \sqrt{2}\end{pmatrix}$
preserving tetrahedral symmetry with edge length $\sqrt{3}$. The parameters of the system (1) remain $\mu=\chi=1$, $\epsilon=10^{-4}$, and $k=10^{-1}$ as in Section 5.1.
Figure 9: Initial tetrahedral density configuration.
We performed simulations for $M_{0}\in\{20,28,40,56,80,112,160\}$ with total simulation time $T=0.4$ and a time step $\tau=10^{-4}$ ($N_{T}=4000$). The grid resolution is fixed at $H=256$. Due to the near blowup scenarios, the smoothness of $\nabla c$ is relatively poor. Therefore, we utilize fourth-order particle-to-grid interpolations and second-order grid-to-particle interpolations to avoid spurious singularities. Particle configurations at $T=0.1$, $0.2$, and $0.4$ are shown in Fig. 10 for three regimes: $M_{0}=20$ (non-blowup), 80 (central blowup), and 160 (two-stage blowup), with other values demonstrating analogous behavior within their respective categories.
Three distinct dynamical regimes are observed. For small $M_{0}$, no blowup phenomenon occurs. At intermediate $M_{0}$, the four clusters merge into a single aggregate prior to a single unified blowup at the center. For large $M_{0}$, the system exhibits a two-stage process: simultaneous blowups initially develop at the original tetrahedral positions, followed by coalescence into a central singularity.
$M_{0}=20,\,T=0.1$
$M_{0}=20,\,T=0.2$
$M_{0}=20,\,T=0.4$
$M_{0}=80,\,T=0.1$
$M_{0}=80,\,T=0.2$
$M_{0}=80,\,T=0.4$
$M_{0}=160,\,T=0.1$
$M_{0}=160,\,T=0.2$
$M_{0}=160,\,T=0.4$
Figure 10: Blowup patterns for different $M_{0}$.
5.4 Collapsing ring blowup dynamics
In the last experiment, we investigate a special blow-up pattern where the mass concentrates on an evolving ring $x_{1}^{2}+x_{2}^{2}=r(t)^{2}$, with the radius $r(t)$ decaying as $t\to T$.
Numerical simulations initialize with a uniform distribution inside the toroidal volume defined by $(1.0-\sqrt{x_{1}^{2}+x_{2}^{2}})^{2}+x_{3}^{2}\leq 0.4^{2}$, with parameters $(k,\epsilon,\mu,\chi)=(10^{-1},10^{-4},1,1)$ in system (1), an initial mass $M_{0}=180$, and a final computational time $T=6\times 10^{-2}$. The spatial domain $\Omega=[-4,4]^{3}$ is discretized with $H=256$ grid points per axis, using $P=2^{20}$ particles, fourth-order particle-to-grid and second-order grid-to-particle interpolations, with a time step $\tau=0.5\times 10^{-4}$, equivalently $N_{T}=1200$.
Figure 11 shows the projected scatter plot of the particle distributions in the $x_{1}$-$x_{2}$ plane at different time steps. The mass progressively collapses into a thin ring with an asymptotically linear density profile as $t\to T$. After complete collapse, numerical stability degrades severely, revealing an extreme sensitivity to asymmetric perturbations. This instability drives particles to aggregate into point-like clusters (singularities) within localized regions, each with a radius of approximately one grid spacing. High particle number and grid resolution are necessary to resolve the ring structure. The original SIPF method ($P=10^{4},H=32$) is not sufficient to resolve the ring structure, as the particles undergo a blow-up at the origin rather than aggregating into a ring.
Figure 11: Toroidal collapse evolution.
6 Conclusions
In this paper, we develop the SIPF-PIC method, a hybrid particle-spectral particle-in-cell method that combines localized particle-grid interpolations with global spectral field solvers accelerated by the FFT method. This new method achieves a computational complexity of $O(P+H^{3}\log H)$, enabling efficient simulations of 3D Keller-Segel (KS) systems with large particle numbers and fine grids. A unified error analysis framework is established to rigorously quantify approximation errors arising from particle-grid interpolations and spectral field evaluations, thus providing convergence guarantees for our SIPF-PIC method in computing non-blow-up solutions of parabolic-parabolic KS systems.
The SIPF-PIC method successfully resolves complex blow-up dynamics beyond classical single-point singularities, such as ring-shaped blow-up patterns in which mass concentrates on an evolving ring structure. Numerical experiments, including convergence tests and 3D visualizations, demonstrate the enhanced precision and scalability of the method in capturing aggregation and near-singular phenomena.
Future work will focus on extending the SIPF-PIC method to other related systems with collective particle-field interactions. Furthermore, we aim to develop adaptive strategies to improve the robustness of the method in handling near-blow-up cases, where high-resolution tracking of density singularities remains a critical challenge.
7 Acknowledgements
ZW was partially supported by NTU SUG-023162-00001 and MOE AcRF Tier 1 Grant RG17/24. JX was partially supported by NSF Grant DMS-2309520, the Swedish Research Council Grant No. 2021-06594 at the Institut Mittag-Leffler in Djursholm, Sweden, and the E. Schrödinger Institute in Vienna, Austria, both during his stay in the fall of 2025. ZZ was partially supported by the National Natural Science Foundation of China (Projects 92470103 and 12171406), the Hong Kong RGC Grant (Projects 17304324 and 17300325), the Seed Funding Programme for Basic Research (HKU), and the Hong Kong RGC Research Fellow Scheme 2025. The computations were performed at the research computing facilities provided by the Greenplanet Cluster at UC Irvine and the Information Technology Services, The University of Hong Kong.